### v2

# Quasi-retracts of groups

Abstract.
A subgroup $H\leq G$ is a quasi-retract of $G$ if there exists a quasi-homomorphism from $G$ to $H$ which restricts to the identity on $H$. In this paper, we study under what conditions a subgroup is a quasi-retract of the whole group. We also show that a subgroup of $F_{2}$ is a nontrivial quasi-retract if and only if it is cyclic. As an intermediate result, we show that for any short exact sequence of groups $1\to H\to G\to G/H\to 1$, the following are equivalent: (i) it is left quasi-split (or equivalently, $H$ is a quasi-retract of $G$). (ii) it is right quasi-split and $H$ almost commutes with a right coset transversal of $H$ in $G$. (iii) $G$ is strictly quasi-isomorphic to $H\times G/H$. This is an analogue to the classical result on split short exact sequence of groups.
Key words and phrases:
Quasi-homomorphisms, extension problem, hyperbolically embedded subgroups, quasi-retracts, short exact sequence
1. Introduction
Let $G$ be a group and $H$ be a group equipped with a proper left-invariant metric $d$ (e.g., a finitely generated group, equipped with a word metric). A map $\phi:G\to H$ is called a quasi-homomorphism if there exists a constant $C>0$ such that
$d(\phi(xy),\phi(x)\phi(y))\leq C$
for all $x,y\in G$. For two maps $\phi,\phi^{\prime}:G\to H$, we define
$\mathrm{dist}(\phi,\phi^{\prime}):=\sup_{x\in G}d(\phi(x),\phi^{\prime}(x)).$
Two quasi-homomorphisms $\phi$ and $\phi^{\prime}$ are called equivalent if $\mathrm{dist}(\phi,\phi^{\prime})<\infty$. In the case when $H$ is discrete (and in this paper we limit ourselves only to this class of groups), $\phi$ is a quasi-homomorphism if and only if the set of defects of $\phi$
$D(\phi)=\{\phi(y)^{-1}\phi(x)^{-1}\phi(xy):x,y\in G\}$
is finite. And two quasi-homomorphisms $\phi,\phi^{\prime}$ are equivalent if and only if $\{\phi(x)^{-1}\phi^{\prime}(x):x\in G\}$ is finite. A quasi-homomorphism with values in $\mathbb{Z}$ (or $\mathbb{R}$, equipped with the standard metric) is called a quasimorphism.
The concept of quasi-homomorphisms goes back to Ulam [13, Chapter 6], who asked if they are close to group homomorphisms. Later, Fujiwara-Kapovich [6] studied this concept to figure out why it is so “hard to construct” quasi-homomorphisms to noncommutative groups which are neither homomorphisms, nor come from quasi-homomorphisms with commutative targets, provided that $H$ is a discrete group. They showed that every quasi-homomorphism is constructible [6, Theorem 1.2]. Roughly speaking, it means that every quasi-homomorphism can be obtained by repeating the following constructions:
  • Lift. A quasi-homomorphism $\phi:G\to H$ is obtained (if possible) by lifting a quasi-homomorphism $\phi:G\to\bar{H}$ where $\bar{H}$ is a quotient group of $H$ with an abelian kernel.
  • Product. A quasi-homomorphism $\phi=(\phi_{1},\ldots,\phi_{n}):G\to\prod_{i=1}^{n}H_{i}$ is obtained by taking product of finitely many quasi-homomorphisms $\phi_{i}:G\to H_{i},1\leq i\leq n$.
  • Composition. A quasi-homomorphism $\phi:G\to H$ is obtained by composing two quasi-homomorphisms $\phi_{1}:G\to K$ and $\phi_{2}:K\to H$.
  • Extension from a finite-index subgroup. A quasi-homomorphism $\phi:G\to H$ is obtained (if possible) by extending a quasi-homomorphism $\phi_{0}:G_{0}\to H$ where $[G:G_{0}]<\infty$.
  • Bounded perturbation. A quasi-homomorphism $\phi:G\to H$ is obtained (if possible) by perturbing a quasi-homomorphism $\phi^{\prime}:G\to H$ a bounded distance, which means that $\mathrm{dist}(\phi,\phi^{\prime})<\infty$.
We note that not all constructions in the above list always give a quasi-homomorphism. A motivation for us to start this work is try to figure out the following question: under what conditions can a quasi-homomorphism from a finite-index subgroup $G_{0}\leq G$ be extended to a quasi-homomorphism from $G$? This question is related to the “exension problem” considered in [8]: Under what conditions can a 1-quasi-cocycle (eg. a quasimorphism) on a subgroup be extended to the whole group? In [8], Hull-Osin solved the problem by assuming the subgroup is hyperbolically embedded. Roughly speaking, hyperbolically embedded subgroups generalize the concept of parabolic subgroups in relatively hyperbolic groups; see [4] for details about hyperbolically embedded subgroups. Moreover, under the same assumption, Frigerio-Pozzetti-Sisto [5] solved a “higher-dimensional extension problem”.
Let $G$ be a group and $H\leq G$ be a hyperbolically embedded subgroup. Observe that if there exists a quasi-homomorphism $r:G\to H$ such that $r|_{H}=Id_{H}$, then any quasimorphism on $H$ can be easily extended to $G$ by composing this quasimorphism with $r$. This gives a generalization to the result of Hull-Osin [8, Theorem 1.4]. Based on this observation, we define a subgroup $H$ of $G$ to be a quasi-retract [^1] This word was first used by Alonso in [1] to denote a certain subordinate relationship between metric spaces. if there exists a quasi-homomorphism $r:G\to H$ such that $r\circ\iota=Id_{H}$ where $\iota:H\to G$ is the inclusion map. This quasi-homomorphism $r$ is called a quasi-retraction from $G$ to $H$. In this paper, we focus on the following questions:
Question 1.
Under what conditions a (finite-index) subgroup is a quasi-retract of the whole group?
Question 2.
Is a hyperbolically embedded subgroup always a quasi-retract?
The first result of this paper is to give a sufficient condition for a finite-index subgroup to be a quasi-retract. For convenience, we say $(G,H)$ is a group pair if $G$ is a group and $H\leq G$ is a subgroup. A core of $H$ in $G$ is defined as $\mathrm{Core}_{G}(H)=\bigcap_{g\in G}g^{-1}Hg$. An equivalent characterization of $\mathrm{Core}_{G}(H)$ is that it is the maximal normal subgroup of $G$ in $H$. For two subsets $H,T$ of a group $G$, we define $\mathcal{C}(H,T)=\{[h,t]=hth^{-1}t^{-1}:h\in H,t\in T\}$. In this case, $H$ and $T$ are commutative (resp. almost commutative) if and only if $\mathcal{C}(H,T)=\{1\}$ (resp. $\mathcal{C}(H,T)$ is finite).
Theorem 1.1.
Let $G$ be a group and $H\leq G$ be a subgroup of finite index. Then the following are equivalent:
  1. (1)
    $\mathrm{Core}_{G}(H)$ is a quasi-retract of $G$.
  2. (2)
    There exists a right coset transversal $T$ containing $1$ of $\mathrm{Core}_{G}(H)$ in $G$ such that $\mathrm{Core}_{G}(H)$ almost commutes with $T$.
Moreover, if one of the above items holds, then $H$ is a quasi-retract of $G$.
In this paper, we denote by $F_{n}$ the free group of rank $n$. The second result of this paper is to determine completely which subgroups of $F_{2}$ are quasi-retracts. As a consequence, we obtain a negative answer to Question 2. A quasi-retract is called nontrivial if it is neither the trivial group nor the whole group.
Theorem 1.2.
A subgroup of $F_{2}$ is a nontrivial quasi-retract if and only if it is a cyclic group. Moreover, there exists a hyperbolically embedded subgroup of $F_{2}$ which is not a quasi-retract.
Another interesting question is that for what kind of finitely generated groups, each of its subgroups is a quasi-retract? We give an answer in the category of elementary groups.
A group is called elementary if it is virtually abelian. An elementary group is called FZ if its center is of finite index. In [10], Nguyen-Wang gave a geometric characterization for FZ groups via a notion of curvature for finitely generated groups. Here, we give an algebraic characterization for FZ groups via quasi-retracts in the category of elementary groups.
Proposition 1.3 (Proposition 3.21).
A finitely generated elementary group $G$ is FZ if and only if every subgroup of $G$ is a quasi-retract.
Another thing worth mentioning is that in [7], Hartnick-Schiweitzer defined a map $\phi:G\to H$ to be a quasi-homomorphism if and only if any quasimorphism on $H$ can be pull-backed to a quasimorphism on $G$ by composing $\phi$. To make a distinction, we shall call such a quasi-homomorphism an HS-quasi-homomorphism. Quasi-homomorphisms and HS-quasi-homomorphisms are not equivalent in general; see [6, Subsection 9.3] for a detailed discussion about different quasi-homomorphisms.
1.1. Tools: Quasi-split short exact sequences
The concept of retraction comes from the classical group theory. A subgroup $H$ is a retract of a group $G$ if there exists a homomorphism $r:G\to H$ such that $r\circ\iota=Id_{H}$ where $\iota:H\to G$ is the inclusion map. This homomorphism $r$ is called a retraction from $G$ to $H$. This notion is widely used in the theory of short exact sequence of groups. Let
(1) $1\to H\to G\to G/H\to 1$
be a short exact sequence of groups. The short exact sequence is left split if there exists a retraction from $G$ to $H$. It is right split if there exists a section $s:G/H\to G$ such that $s$ is a homomorphism. Similarly, we say the short exact sequence is left quasi-split if there exists a quasi-retraction from $G$ to $H$. It is right quasi-split if there exists a section $s:G/H\to G$ such that $s$ is a quasi-homomorphism. We remark that the notion of right quasi-split here is the same as the notion of quasisplit in [6, Subsection 2.4.1].
Theorem 1.4 (Classical version).
[…]
  1. (1)
    The short exact sequence is left split. Or equivalently, $H$ is a retract of $G$.
  2. (2)
    There exists a section $s:G/H\to G$ such that $s$ is a homomorphism and $H$ commutes with $s(G/H)$.
  3. (3)
    There is an isomorphism $\phi:G\to H\times G/H$ such that the diagram commutes:
[…]
    where the bottom row is the short exact sequence for a direct product.
A section $s$ is normalized if $s(1)=1$. A quasi-homomorphism $\phi:G\to H$ is a quasi-isomorphism if it admits a quasi-inverse, i.e., a quasi-homomorphism $\phi^{\prime}:H\to G$ such that $\mathrm{dist}(\phi^{\prime}\circ\phi,Id_{G})<\infty,\mathrm{dist}(\phi\circ\phi^{\prime},Id_{H})<\infty$. A quasi-isomorphism is strict if $\phi^{\prime}=\phi^{-1}$. With these notations, we have the following result which is an analogue to Theorem 1.4.
Theorem 1.5 (Theorem 3.15).
[…]
  1. (1)
    The short exact sequence is left quasi-split. Or equivalently, $H$ is a quasi-retract of $G$.
  2. (2)
    There exists a normalized section $s:G/H\to G$ such that $s$ is a quasi-homomorphism and $H$ almost commutes with $s(G/H)$.
  3. (3)
    There is a strict quasi-isomorphism $\phi:G\to H\times G/H$ such that the diagram commutes:
[…]
    where the bottom row is the short exact sequence for a direct product.
When $H$ is a finite-index normal subgroup of $G$, any normalized section $s:G/H\to G$ is a quasi-homomorphism since $G/H$ is finite. Thus as a corollary of Theorem 1.5, we have that
Corollary 1.6.
Let $H$ be a finite-index normal subgroup of $G$. Then $H$ is a quasi-retract of $G$ if and only if there exists a right coset transversal $T$ containing $1$ of $H$ in $G$ such that $H$ almost commutes with $T$. Moreover, if one of the above conditions holds, then $H$ is quasi-isomorphic to $G$.
This corollary, together with Lemma 3.18 , completes the proof of Theorem 1.1.
Structure of the paper
The paper is organized as follows. Section 2 is devoted to defining some notations and studying some elementary properties of quasi-homomorphisms. In Subsection 3.1, we first give some examples and non-examples of quasi-retractions and then complete the proof of Theorem 1.2. In Subsection 3.2, we introduce the classical theory of split short exact sequence and prove Theorem 1.5 first. By utilizing Theorem 1.5 and Lemma 3.18, we complete the proof of Theorem 1.1 and then Proposition 1.3.
Acknowledgments
We are grateful to Zhenguo Huangfu and Bingxue Tao for many helpful suggestions on the first draft. R. W. is supported by NSFC No.12471065 & 12326601 and in part by Science and Technology Commission of Shanghai Municipality (No. 22DZ2229014).
2. Preliminaries
2.1. Definitions and notations
Let $H$ be a group with a proper left-invariant metric $d$. For a subset $A\subset H$ and a constant $C>0$, we denote $\mathcal{N}_{C}(A):=\{h\in H:d(h,A)\leq C\}$.
Let $S\subset H$ be a finite subset. For convenience, we will always write
$h\sim_{S}h^{\prime}$
if $h=h^{\prime}s$ with $h,h^{\prime}\in H,s\in S$. For example, for a quasi-homomorphism $\phi:G\to H$ with $D=D(\phi)$, by the definition,
$\phi(ab)\sim_{D}\phi(a)\phi(b)\text{ and }\phi(abc)\sim_{D}\phi(a)\phi(bc)\sim_{D}\phi(a)\phi(b)\phi(c)$
for $a,b,c\in G$.
Definition 2.1.
[6, Definition 2.1] Suppose that a map $\phi:G\to H$ between groups has the property that $\phi(G)$ is contained in a subgroup $J<H$, $J$ contains a finite normal subgroup $K\lhd J$, such that the projection $\bar{\phi}:G\to\bar{J}=J/K$ is a homomorphism. We then will refer to $\phi$ as an almost homomorphism, it is a homomorphism modulo a finite normal subgroup (in the range of $\phi$).
Clearly, every almost homomorphism is a quasi-homomorphism.
For a subset $D$ of a group $H$ and $n\geq 2$ we will use the notation $D^{n}$ to denote the subset of $H$ consisting of products of at most $n$ elements of $D$. More generally, for two subsets $A,B\subset H$ we let
$A\cdot B=\{ab:a\in A,b\in B\}.$
We will use the notation $D^{-1}$ for the set of inverses of elements of $D$. Then
$h\sim_{D}h^{\prime}\Longleftrightarrow h^{\prime}\sim_{D^{-1}}h.$
Given a group $H$ and its subgroup $A$ we let $N_{H}(A)$ and $Z_{H}(A)$ denote the normalizer and the centralizer of $A$ in $H$ respectively. We also denote $\langle\langle A\rangle\rangle$ as the normal closure of $A$, i.e., the smallest normal subgroup containing $A$ in $H$.
Given a group $G$ with a finite generating set $S$, we denote by $d_{S}$ the word metric on $G$.
Given a quasi-homomorphism $\phi:G\to H$, we denote by $\Delta_{\phi}$ the defect subgroup of $H$ which is finitely generated by $D(\phi)$.
2.2. Elementary properties of quasi-homomorphisms
We collect some basic facts about quasi-homomorphisms. See [6, Subsections 2.2, 2.3, 3.1] for their proofs.
Lemma 2.2.
Let $\phi:G\to H$ be a quasi-homomorphism with $D=D(\phi)$.
  1. (1)
    For any $x\in G$, $\phi(x)^{-1}\sim_{D^{2}}\phi(x^{-1})$.
  2. (2)
    For any $h\in\phi(G)$, $h^{-1}Dh\subset D^{2}D^{-1}$.
  3. (3)
    $\phi(G)\cup\phi(G)^{-1}\subset N_{H}(\Delta_{\phi})$ and the induced map $\bar{\phi}:G\to N_{H}(\Delta_{\phi})/\Delta_{\phi}$ is a homomorphism.
  4. (4)
    There exists a constant $C$ such that $\phi(G)\subset\mathcal{N}_{C}(Z_{H}(\Delta_{\phi}))$.
For a map $\phi:G\to H$, we define the dual defect set of $\phi$ as
$\tilde{D}(\phi)=\{\phi(x)\phi(y)\phi(xy)^{-1}:x,y\in G\}.$
Lemma 2.3.
A map $\phi:G\to H$ is a quasi-homomorphism if and only if $\tilde{D}(\phi)$ is finite.
Proof.
By definition of quasi-homomorphisms, it is equivalent to show that $D(\phi)$ is finite if and only if $\tilde{D}(\phi)$ is finite. By symmetry, we only need to show that $\tilde{D}(\phi)$ is finite when $D(\phi)$ is finite.
Denote $D=D(\phi)$. Lemma 2.2 (2) shows that $h^{-1}Dh\subset D^{2}D^{-1}$ for any $h\in\phi(G)$. Thus, $h^{-1}D^{-1}h\subset DD^{-2}$. Moreover, any $h\in\phi(G)$ can be written as $h=\phi(x^{-1})$ for some $x^{-1}\in G$. By Lemma 2.2 (1), $\phi(x^{-1})\sim_{D^{-2}}\phi(x)^{-1}$. Together with the above discussions, one has that
$hD^{-1}h^{-1}=\phi(x^{-1})D^{-1}\phi(x^{-1})^{-1}\subset\phi(x)^{-1}D^{-3}D^{2}\phi(x)\subset D^{\prime}$
where $D^{\prime}=(DD^{-2})^{3}(D^{2}D^{-1})^{2}$.
For any $x,y\in G$, there exists $s\in D$ such that $\phi(xy)=\phi(x)\phi(y)s$. Therefore,
$\phi(x)\phi(y)\phi(xy)^{-1}=\phi(x)\phi(y)s^{-1}\phi(y)^{-1}\phi(x)^{-1}\in\phi(x)D^{\prime}\phi(x)^{-1}\subset D^{\prime\prime}$
where $D^{\prime\prime}=(D^{\prime-1}D^{\prime 2})^{3}(D^{\prime-2}D^{\prime})^{2}$ is a finite set. Since $x,y\in G$ are arbitrary, the above formula shows that $\tilde{D}(\phi)\subset D^{\prime\prime}$, which completes the proof. ∎
Let $(G,H)$ be a group pair and $T$ be a right coset transversal of $H$ in $G$. It is clear that any element $g\in G$ can be written uniquely as $g=g_{H}g_{T}$ where $g_{H}\in H$ and $g_{T}\in T$. The following simple lemma will be used frequently in later proofs.
Lemma 2.4.
Let $(G,H)$ be a group pair and $T$ be a right coset transversal of $H$ in $G$. If $H$ is a normal subgroup of $G$, then the product of any two elements $g,g^{\prime}\in G$ satisfies that $(gg^{\prime})_{H}=g_{H}(g_{T}g^{\prime}_{H}g_{T}^{-1})(g_{T}g^{\prime}_{T})_{H}$ and $(gg^{\prime})_{T}=(g_{T}g^{\prime}_{T})_{T}$. Moreover, $(g_{T}g^{\prime}_{T})_{H}\in\tilde{D}(s)$ where $s:G/H\to G$ is a section such that $T=s(G/H)$.
Proof.
It is straightforward to calculate that
$gg^{\prime}=g_{H}g_{T}g^{\prime}_{H}g^{\prime}_{T}=g_{H}(g_{T}g^{\prime}_{H}g_{T}^{-1})g_{T}g^{\prime}_{T}=g_{H}(g_{T}g^{\prime}_{H}g_{T}^{-1})(g_{T}g^{\prime}_{T})_{H}(g_{T}g^{\prime}_{T})_{T}.$
Since $H\lhd G$, $g_{T}g^{\prime}_{H}g_{T}^{-1}\in H$. By the uniqueness of decomposition, one gets that $(gg^{\prime})_{H}=g_{H}(g_{T}g^{\prime}_{H}g_{T}^{-1})(g_{T}g^{\prime}_{T})_{H}$ and $(gg^{\prime})_{T}=(g_{T}g^{\prime}_{T})_{T}$. Moreover, let $\pi:G\to G/H$ be the quotient map and $s:G/H\to G$ be a section such that $T=s(G/H)$. This means that for any $g=g_{H}g_{T}\in G$, $s(\pi(g))=s(\pi(g_{T}))=g_{T}$. Then $s(\pi(gg^{\prime}))=(gg^{\prime})_{T}=(g_{T}g^{\prime}_{T})_{T}$. Finally, it follows from $g_{T}g^{\prime}_{T}=(g_{T}g^{\prime}_{T})_{H}(g_{T}g^{\prime}_{T})_{T}$ that
$(g_{T}g^{\prime}_{T})_{H}=g_{T}g^{\prime}_{T}(g_{T}g^{\prime}_{T})_{T}^{-1}=s(\pi(g))s(\pi(g^{\prime}))s(\pi(gg^{\prime}))^{-1}\in\tilde{D}(s).$
Recall that a quasi-homomorphism $\phi:G\to H$ is a quasi-isomorphism if there exists a quasi-homomorphism $\phi^{\prime}:H\to G$, called quasi-inverse of $\phi$ such that $\mathrm{dist}(\phi^{\prime}\circ\phi,Id_{G})<\infty$ and $\mathrm{dist}(\phi\circ\phi^{\prime},Id_{H})<\infty$. Moreover, a quasi-isomorphism is called strict if $\phi^{\prime}=\phi^{-1}$.
Lemma 2.5.
Quasi-isomorphism and strict quasi-isomorphism are two equivalent relations on discrete groups.
Proof.
Both reflexivity and symmetry are easy to see. We only need to verify the transitivity. Let $\phi_{1}:G\to H$ and $\phi_{2}:H\to K$ be two quasi-isomorphisms between discrete groups. By definition, there exist two quasi-homomorphisms $\phi_{1}^{\prime}:H\to G$ and $\phi_{2}^{\prime}:K\to H$ and four finite subsets $A_{1}\subset G,B_{1},A_{2}\subset H,B_{2}\subset K$ such that for any $g\in G,h\in H,k\in K$,
$\phi_{1}^{\prime}(\phi_{1}(g))\sim_{A_{1}}g,\quad\phi_{1}(\phi_{1}^{\prime}(h))\sim_{B_{1}}h,\quad\phi_{2}^{\prime}(\phi_{2}(h))\sim_{A_{2}}h,\quad\phi_{2}(\phi_{2}^{\prime}(k))\sim_{B_{2}}k.$
It follows that
$\phi_{2}^{\prime}\circ\phi_{2}(\phi_{1}(g))\sim_{A_{2}}\phi_{1}(g),\quad\phi_{1}\circ\phi_{1}^{\prime}(\phi_{2}^{\prime}(k))\sim_{B_{1}}\phi_{2}^{\prime}(k).$
By composing $\phi_{1}^{\prime}$ or $\phi_{2}$ on the left, we get
$\phi_{1}^{\prime}\circ\phi_{2}^{\prime}\circ\phi_{2}(\phi_{1}(g))\sim_{\phi_{1}^{\prime}(A_{2})D(\phi_{1}^{\prime})}\phi_{1}^{\prime}(\phi_{1}(g))\sim_{A_{1}}g$
and
$\phi_{2}\circ\phi_{1}\circ\phi_{1}^{\prime}(\phi_{2}^{\prime}(k))\sim_{\phi_{2}(B_{1})D(\phi_{2})}\phi_{2}(\phi_{2}^{\prime}(k))\sim_{B_{2}}k.$
Denote $\phi=\phi_{2}\circ\phi_{1}:G\to K$ and $\phi^{\prime}=\phi_{1}^{\prime}\circ\phi_{2}^{\prime}:K\to G$ as two composite maps. They are both quasi-homomorphisms since a composition of quasi-homomorphisms is still a quasi-homomorphism. The above formulas show that both $\{g^{-1}\phi^{\prime}(\phi(g)):g\in G\}$ and $\{k^{-1}\phi(\phi^{\prime}(k)):k\in K\}$ are finite. Therefore, $\phi,\phi^{\prime}$ are quasi-isomorphisms, which shows that quasi-isomorphism is an equivalent relation on discrete groups. The same proof shows that strict quasi-isomorphism is also an equivalent relation on discrete groups. ∎
3. Quasi-retraction
Definition 3.1.
Let $(G,H)$ be a group pair. Classically a group homomorphism $r:G\to H$ is called a retraction if it is left-inverse to the inclusion $\iota:H\to G$. We thus call a quasi-homomorphism $r:G\to H$ a quasi-retraction if $r\circ\iota=Id_{H}$. If such a map exists, then $H$ is called a $quasi-retract$ of G.
At first, we give some basic properties of quasi-retractions.
Lemma 3.2.
Let $\phi:G\to H$ be a surjective quasi-homomorphism. Then $\Delta_{\phi}$ is a virtually abelian normal subgroup of $H$.
Proof.
Since $\phi$ is surjective, Lemma 2.2 (3) shows that $H=\phi(G)=N_{H}(\Delta_{\phi})$. Thus $\Delta_{\phi}\lhd H$. By Lemma 2.2 (4), there exists a constant $C$ such that $\phi(G)=H\subset\mathcal{N}_{C}(Z_{H}(\Delta_{\phi}))$. This implies that $[H:Z_{H}(\Delta_{\phi})]<\infty$. Denote $\Delta_{\phi}^{\prime}=\Delta_{\phi}\cap Z_{H}(\Delta_{\phi})$. Note that $\Delta_{\phi}^{\prime}$ is an abelian group. According to the isomorphism theorem of groups, $[\Delta_{\phi}:\Delta_{\phi}^{\prime}]\leq[H:Z_{H}(\Delta_{\phi})]<\infty$, which implies that $\Delta_{\phi}$ is virtually abelian. ∎
Lemma 3.3.
Let $r:G\to H$ be a quasi-retraction with $D=D(r)$. Denote $\tilde{\Delta}_{r}$ as the normal closure of $\Delta_{r}$ in $G$, i.e., $\tilde{\Delta}_{r}=\langle\langle\Delta_{r}\rangle\rangle$.
  1. (1)
    Foe any $g\in G$, $r(g)^{-1}g\in r^{-1}(D)$.
  2. (2)
    $r(\tilde{\Delta}_{r})\subset\Delta_{r}$.
  3. (3)
    If $G$ is finitely generated, then so is $H$.
  4. (4)
    Suppose that $G$ is generated by a finite set $S$. Then $r$ is a Lipchitz map with respect to $d_{S}$ on $G$ and any proper word metric on $H$.
Proof.
(1) For any $g\in G$, $r(g)^{-1}\in H$. Then $r(r(g)^{-1}g)\sim_{D}r(r(g)^{-1})r(g)=r(g)^{-1}r(g)=1$.
(2) Since $r$ is a surjective map, Lemma 3.2 shows that $\Delta_{r}$ is a normal subgroup of $H$. For any $g\in G$, Lemma 2.2 (1) shows that $r(g^{-1})\sim_{D^{-2}}r(g)^{-1}$. Then for any $s\in\Delta_{r}$, one has that
$r(gsg^{-1})\sim_{D^{2}}r(g)r(s)r(g^{-1})\sim_{D^{-2}}r(g)sr(g)^{-1}\in\Delta_{r}.$
Denote $S:=\{gsg^{-1}:g\in G,s\in\Delta_{r}\}$. The above formula shows that $r(S)\subset\Delta_{r}$. Since $\tilde{\Delta}_{r}$ is generated by $S$, any element $g\in\tilde{\Delta}_{r}$ can be written as $g=s_{1}\cdots s_{n}$ where $s_{i}\in S,1\leq i\leq n$. Therefore,
$r(g)=r(s_{1}\cdots s_{n})\sim_{D^{n-1}}r(s_{1})\cdots r(s_{n})\in r(S)^{n}\subset\Delta_{r}.$
(3) Let $S$ be a finite generating set of $G$. For any $h\in H\leq G$, we denote $h=s_{1}\cdots s_{n}$ where $s_{i}\in S,1\leq i\leq n$. Since $r|_{H}=Id_{H}$,
$h=r(h)=r(s_{1}\cdots s_{n})\sim_{D^{n-1}}r(s_{1})\cdots r(s_{n})\subset r(S)^{n}.$
This implies that $r(S)\cup D\subset H$ is a finite generating set of $H$.
(4) By Item (3), $S^{\prime}=r(S)\cup D$ is a finite generating set of $H$. Fix arbitrary two elements $g,g^{\prime}\in G$. Denote $n=d_{S}(g,g^{\prime})$. Namely, $g^{-1}g^{\prime}=s_{1}\cdots s_{n}$ where $s_{i}\in S,1\leq i\leq n$. Then
$r(g^{-1}g^{\prime})=r(s_{1}\cdots s_{n})\sim_{D^{n-1}}r(s_{1})\cdots r(s_{n})$
which implies that $d_{S^{\prime}}(1,r(g^{-1}g^{\prime}))\leq 2n-1$. Moreover, there exists $a\in D$ such that
$d_{S^{\prime}}(r(g),r(g^{\prime}))=d_{S^{\prime}}(r(g),r(g)r(g^{-1}g^{\prime})a)=d_{S^{\prime}}(1,r(g^{-1}g^{\prime})a)\leq 2n-1+1=2d_{S}(g,g^{\prime}).$
Since $g,g^{\prime}\in G$ are arbitrary, we get that $r$ is a Lipchitz map. ∎
Remark 3.4.
By [1, Theorem 8], together with Lemma 3.3 (4), one has that $H$ is of type $F_{n}$ (or $FP_{n}$) whenever $G$ is of type $F_{n}$ (or $FP_{n}$). This is a generalization of Lemma 3.3 (3). By the way, a hyperbolically embedded subgroup shares the same conclusion [4, Theorem 2.11].
3.1. Examples and non-examples of quasi-retractions
Next, we collect some examples and non-examples of quasi-retractions.
Definition 3.5.
[12] A finitely generated group $G$ has property (PH) if $G$ contains a subgroup $G_{0}$ of finite index satisfying that there exist finitely many Gromov-hyperbolic spaces $X_{1},\ldots,X_{n}$ on which $G_{0}$ acts coboundedly and the induced diagonal action of $G_{0}$ on the product $\prod_{i=1}^{n}X_{i}$ with $\ell^{1}$-metric is proper.
By definition, any finitely generated virtually abelian group has property (PH).
By [6, Lemma 2.9], a central extension of groups $1\to H\to G\to G/H\to 1$ has bounded Euler class if and only if the central extension is right quasi-split, i.e. there exists a section $s:G/H\to G$ such that $s$ is a quasi-homomorphism.
Lemma 3.6.
Let $(G,H)$ be a group pair. Then $H$ is a quasi-retract of $G$ whenever the pair satisfies one of the following:
  1. (1)
    $G$ is an arbitrary group and $H$ is a finite subgroup.
  2. (2)
    $G$ is a direct product of a collection of groups and $H$ is some direct factor.
  3. (3)
    $G$ is a free product of a collection of groups and $H$ is some free factor.
  4. (4)
    $G$ has property (PH) and $H\leq Z(G)$.
  5. (5)
    $G$ is a finitely generated abelian group and $H$ is an arbitrary subgroup.
  6. (6)
    $H\leq Z(G)$ and the central extension $1\to H\to G\to G/H\to 1$ has bounded Euler class.
Proof.
(1) is clear by mapping every element in $H$ to itself and every element outside $H$ to $1$. (2) and (3) are clear since there exists a natural epimorphism from $G$ to $H$. (6) is given by [12, Lemma 5.1]. By [12, Theorem 1.5], (4) is transformed into (6). (5) follows directly from (4). ∎
We also collect some non-examples of quasi-retractions.
Recall that a quasi-homomorphism with values in $\mathbb{R}$ equipped with the standard metric is called a quasimorphism. A quasimorphism $\phi:G\to\mathbb{R}$ is called homogeneous if $\phi(g^{n})=n\phi(g)$ for all $g\in G,n\in\mathbb{Z}$. According to [3, Lemma 2.21], any quasimorphism $\phi:G\to\mathbb{R}$ has a homogenization which is a homogeneous quasimorphism $\bar{\phi}:G\to\mathbb{R}$ defined by $\bar{\phi}(g)=\lim_{n\to\infty}\phi(g^{n})/n$.
Example 3.7.
  1. (1)
    Let $G=\langle t,s\mid s^{2}=(st)^{2}=1\rangle$ be an infinite dihedral group and $H=\langle t\rangle$. Clearly the map $\eta:H\to\mathbb{R},t^{n}\mapsto n$ gives a homomorphism from $H$ to $\mathbb{R}$. If $\phi:G\to H$ is a quasi-retraction, then the homogenization $\bar{\varphi}$ of the composite map $\varphi=\eta\circ\phi$ is a nonzero homogeneous quasimorphism on both $G$ and $H$. Since a homogeneous quasimorphism takes constant values in the same conjugacy class, $\bar{\varphi}(t)=\bar{\varphi}(sts^{-1})=\bar{\varphi}(t^{-1})=-\bar{\varphi}(t)$, which forces that $\bar{\varphi}(t)=0$. This contradicts with $\bar{\varphi}|_{H}\neq 0$. Therefore, $H$ is not a quasi-retract of $G$.
  2. (2)
    Let $G=\langle a,b,c\mid[a,c]=[b,c]=1,[a,b]=c\rangle$ be the Heisenberg group and $H=\langle c\rangle=Z(G)$. Since the central extension $1\to H\to G\to G/H\to 1$ is not right quasi-split [6, Lemma 2.9, Example 2.11], as a result of Theorem 1.5, $H$ is not a quasi-retract of $G$. However, $G$ quasi-retracts to $\langle a\rangle$ since $G$ has a homomorphism to $\langle a\rangle$ by sending $a,b,c$ to $a,1,1$ respectively.
  3. (3)
    Let $G=BS(1,n)=\langle a,t\mid tat^{-1}=a^{n}\rangle,n\neq 0,1$ be a Baumslag-Solitar group and $H=\langle a\rangle$. Suppose that there exists a quasi-retraction $r:G\to H$ with a defect set $D=D(r)$. Then one has
    $a^{mn}=r(a^{mn})=r(ta^{m}t^{-1})\sim_{D^{2}}r(t)r(a^{m})r(t^{-1})=r(a^{m})r(t)r(t^{-1})\sim_{r(t)r(t^{-1})}r(a^{m})=a^{m}$
    for any $m\in\mathbb{N}$. Since $n\neq 0,1$, $|m(n-1)|\to\infty$ as $m\to\infty$. This is a contradiction to the above formula. Hence, $H$ is not a quasi-retract of $G$. However, $G$ quasi-retracts to $\langle t\rangle$ since $G$ has a homomorphism to $\langle t\rangle$ by sending $t,a$ to $t,1$ respectively.
In the end of this subsection, we are going to give a characterization of quasi-retracts of $F_{2}$ and prove Theorem 1.2. Recall that a quasi-retract is called nontrivial if it is not the trivial group or the whole group. A group element $g\in G$ is called primitive if $g$ can not be represented as a nonzero power of any $h\in G\setminus\{g,g^{-1}\}$. A group $G$ is called Hopfian if any surjective homomorphism from $G$ to $G$ is an isomorphism.
At first, the following result gives a characterization of retracts (NOT quasi-retracts) of $F_{2}$.
Lemma 3.8.
Let $G=F_{2}$. Any nontrivial retract of $G$ is a cyclic subgroup generated by a primitive element.
Proof.
Let $H\leq G$ be a nontrivial retract of $G$. That is, there exists a homomorphism $r:G\to H$ such that $r|_{H}=Id$. Since $r$ is a surjective homomorphism, $\mathrm{rk}(H)\leq\mathrm{rk}(G)=2$. Since $H$ is nontrivial, $\mathrm{rk}(H)\geq 1$.
If $\mathrm{rk}(H)=2$, then $H\cong F_{2}$ since any subgroup of a free group is also free. Another well-known fact is that any free group of finite rank is Hopfian [11, Theorem 3.3]. This implies that $r$ is an isomorphism, which contradicts with that $H\neq G$. Therefore $\mathrm{rk}(H)=1$.
Let $H=\langle g\rangle$. Suppose that $g$ is not primitive, i.e. $g=h^{n}$ for some $h\neq 1\in G,n\geq 2$. Denote $r(h)=h^{k}$ for some $k\in\mathbb{Z}$. Then $h^{n}=g=r(g)=r(h^{n})=r(h)^{n}=h^{kn}$ which implies that $k=1$. However, $r(h)=h\notin H$ which is impossible. Therefore, $g$ is a primitive element. ∎
Now, we can characterize quasi-retracts of $F_{2}$.
Lemma 3.9.
Let $G=F_{2}$ and $H$ be a subgroup of $G$. Then $H$ is a nontrivial quasi-retract of $G$ if and only if $H$ is a cyclic subgroup.
Proof.
$\Leftarrow$”: Let $H=\langle w\rangle$. Up to conjugation, we can assume that $w$ is a cyclically reduced word in $G$. Let $h_{w}$ be the corresponding Brooks counting quasimorphism, which is a quasimorphism on $G$ satisfying $h_{w}(w^{n})=n$ for $n\in\mathbb{Z}$; see [3, Subsection 2.3.3] for a reference. Then the map $r:G\to H$ defined by $r(g)=w^{h_{w}(g)}$ is a quasi-retract.
$\Rightarrow$”: Let $H\leq G$ be a nontrivial quasi-retract of $G$ and $r:G\to H$ be a quasi-retraction. By Lemma 3.3 (3), $H$ is finitely generated. By Lemma 3.2, $\Delta_{r}$ is a virtually abelian normal subgroup of $H$. Recall from the definition of defect subgroups that $\Delta_{r}$ is finitely generated. Since a subgroup of a free group is also free, $H\cong F_{n}$ for some $n\geq 1$ and $\Delta_{r}$ is cyclic. If $\Delta_{r}=\{1\}$, then $r$ is a retraction and the remaining proof is completed by Lemma 3.8. If $\Delta_{r}\neq\{1\}$, then $H$ is virtually cyclic since a nontrivial finitely generated normal subgroup of a free group must be of finite index. Moreover, since $H$ is free, one gets that $H$ is also cyclic. ∎
A subgroup $H\leq G$ is called malnormal if $H\cap gHg^{-1}=\{1\}$ for any $g\in G\setminus H$.
Proposition 3.10.
There exists a hyperbolically embedded subgroup of $F_{2}$ which is not a quasi-retract.
Proof.
In [2, Example 1], Baumslag-Miasnikov-Remeslennikov constructed an example of a finitely generated malnormal subgroup $H$ of $F_{2}$ which has rank greater than 2. Hence, Lemma 3.9 shows that it can not be a quasi-retract of $F_{2}$. Since any finitely generated subgroup of a free group is quasi-convex, as a result of [4, Theorem 2.7], $H$ is a hyperbolically embedded subgroup of $F_{2}$. ∎
By combining Proposition 3.10 with Lemma 3.9, we complete the proof of Theorem 1.2.
3.2. Quasi-split short exact sequence
Recall from the Introduction that, a short exact sequence $1\to H\to G\to G/H\to 1$ is left split if $H$ is a retract of $G$ and right split if it admits a section $s:G/H\to G$ that is a homomorphism. Comparing with the classical notion, we say the short exact sequence is left quasi-split if $H$ is a quasi-retract of $G$ and right quasi-split if it admits a section $s:G/H\to G$ that is a quasi-homomorphism. One simple observation is that whenever $H$ or $G/H$ is finite, the above short exact sequence is always right quasi-split since the defect set of any section is a finite subset of $H$.
Recall in the classical group theory, we have the following results.
Theorem 3.11 (Classical version).
Let $1\to H\to G\to G/H\to 1$ be a short exact sequence of groups. The following are equivalent:
  1. (1)
    The short exact sequence $1\to H\to G\to G/H\to 1$ is left split. Or equivalently, $H$ is a retract of $G$.
  2. (2)
    There exists a section $s:G/H\to G$ such that $s$ is a homomorphism and $H$ commutes with $s(G/H)$.
  3. (3)
    There is an isomorphism $\phi:G\to H\times G/H$ such that the diagram commutes:
[…]
    where the bottom row is the short exact sequence for a direct product.
Before giving the analogue of Theorem 3.11, we need the following two lemmas. Recall that for any two subsets $H,T$ of a group $G$, we denote $\mathcal{C}(H,T):=\{[h,t]=hth^{-1}t^{-1}:h\in H,t\in T\}$.
Lemma 3.12.
Let $G$ be a group and $H\lhd G$ be a normal subgroup. For any subset $T\subset G$, $\mathcal{C}(H,T)=\mathcal{C}(T^{-1},H)=\mathcal{C}(H,T^{-1})^{-1}$.
Proof.
Let $T\subset G$ be a subset. Since $[t^{-1},h]=t^{-1}hth^{-1}=(ht^{-1}h^{-1}t)^{-1}=[h,t^{-1}]^{-1}$ for all $h\in H,t\in T$, it is clear that $\mathcal{C}(T^{-1},H)=\mathcal{C}(H,T^{-1})^{-1}$. Thus we only need to verify that $\mathcal{C}(H,T)=\mathcal{C}(T^{-1},H)$. Since $H\lhd G$, $tht^{-1},t^{-1}ht\in H$ for any $h\in H,t\in T$. Denote $h^{t}:=tht^{-1}$ and $h^{t^{-1}}=t^{-1}ht$. Then $(h^{-1})^{t}=th^{-1}t^{-1}=(h^{t})^{-1}$ and $(h^{-1})^{t^{-1}}=t^{-1}h^{-1}t=(h^{t^{-1}})^{-1}$. Therefore,
$[h,t]=hth^{-1}t^{-1}=t^{-1}(tht^{-1})t(th^{-1}t^{-1})=t^{-1}h^{t}t(h^{-1})^{t}=[t^{-1},h^{t}]\in\mathcal{C}(T^{-1},H)$
and similarly
$[t^{-1},h]=t^{-1}hth^{-1}=(t^{-1}ht)t(t^{-1}h^{-1}t)t^{-1}=h^{t^{-1}}t(h^{-1})^{t^{-1}}t^{-1}=[h^{t^{-1}},t]\in\mathcal{C}(H,T).$
Since $h\in H,t\in T$ are arbitrary, the above formulas show that $\mathcal{C}(H,T)\subset\mathcal{C}(T^{-1},H)\subset\mathcal{C}(H,T)$. Thus we get $\mathcal{C}(H,T)=\mathcal{C}(T^{-1},H)$ and the conclusion follows. ∎
The following lemma gives a useful necessary condition when a subgroup is a quasi-retract.
Lemma 3.13.
Let $(G,H)$ be a group pair. If $H$ is a quasi-retract of $G$, then for any normal subgroup $H_{0}\lhd G$ in $H$, there exists a right coset transversal $T$ containing $1$ of $H$ in $G$ such that $\mathcal{C}(H_{0},T)$ is finite.
Proof.
By assumption, there exists a quasi-retraction $r$ from $G$ to $H$. Denote by $D=D(r)$ the finite defect set of $r$. Since $r(1)=1=r(gg^{-1})\sim_{D}r(g)r(g^{-1})$, one has that $r(g^{-1})\sim_{D^{-1}}r(g)^{-1}$ for any $g\in G$. Let $T$ be an arbitrary right coset transversal containing $1$ of $H$ in $G$. For each $t\in T$, we denote $t^{\prime}:=r(t)^{-1}t$ and $T^{\prime}=\{t^{\prime}:t\in T\}$. Clearly $T^{\prime}$ gives another right coset transversal containing $1$ of $H$ in $G$. And Lemma 3.3 (1) shows that $r(T^{\prime})\subset D$.
Since $H_{0}\lhd G$ and $r|_{H_{0}}=Id_{H_{0}}$, one has that for any $h\in H_{0},t^{\prime}\in T^{\prime}$,
$t^{\prime}ht^{\prime-1}=r(t^{\prime}ht^{\prime-1})\sim_{D^{2}}r(t^{\prime})r(h)r(t^{\prime-1})=r(t^{\prime})hr(t^{\prime-1})\sim_{D^{-1}}r(t^{\prime})hr(t^{\prime})^{-1}\sim_{D^{-1}}r(t^{\prime})h.$
By multiplying $h^{-1}r(t^{\prime})^{-1}$ to the left, the above formula transforms into
$h^{-1}r(t^{\prime})^{-1}t^{\prime}ht^{\prime-1}\sim_{D^{-2}D^{2}}1.$
Denote $\tilde{t}:=r(t^{\prime})^{-1}t^{\prime}$. Therefore, we get
$[h^{-1},\tilde{t}]=h^{-1}\tilde{t}h\tilde{t}^{-1}=h^{-1}r(t^{\prime})^{-1}t^{\prime}ht^{\prime-1}r(t^{\prime})\sim_{D}h^{-1}r(t^{\prime})^{-1}t^{\prime}ht^{\prime-1}\sim_{D^{-2}D^{2}}1.$
Since $h\in H_{0}$ is arbitrary and $\tilde{T}=\{\tilde{t}:t\in T\}$ is still a right coset transversal containing $1$ of $H$ in $G$, the conclusion follows. ∎
Remark 3.14.
We remark that the inverse of Lemma 3.13 is not true. For example, let $G$ be a Heisenberg group and $H$ be its center. Then $\mathcal{C}(H_{0},T)=\{1\}$ since $H_{0}\leq H=Z(G)$. However, Example 3.7 (2) shows that $H$ is not a quasi-retract of $G$.
A section is called normalized if $s(1)=1$.
Theorem 3.15.
[…]
  1. (1)
    The short exact sequence $1\to H\to G\to G/H\to 1$ is left quasi-split. Or equivalently, $H$ is a quasi-retract of $G$.
  2. (2)
    There exists a normalized section $s:G/H\to G$ such that $s$ is a quasi-homomorphism and $H$ almost commutes with $s(G/H)$: $\mathcal{C}(H,s(G/H))$ is finite.
  3. (3)
    There is a strict quasi-isomorphism $\phi:G\to H\times G/H$ such that the diagram commutes:
[…]
    where the bottom row is the short exact sequence for a direct product.
Proof.
[…]
For any $g=g_{H}g_{T},g^{\prime}=g^{\prime}_{H}g^{\prime}_{T}\in G$, one has that $gg^{\prime}=g_{H}g_{T}g^{\prime}_{H}g^{\prime}_{T}=(gg^{\prime})_{H}(gg^{\prime})_{T}$. Therefore,
$r(g^{\prime})^{-1}r(g)^{-1}r(gg^{\prime})={g^{\prime}_{H}}^{-1}{g_{H}}^{-1}(gg^{\prime})_{H}={g^{\prime}_{H}}^{-1}g_{T}g^{\prime}_{H}g^{\prime}_{T}{(gg^{\prime})_{T}}^{-1}=[{g^{\prime}_{H}}^{-1},g_{T}]g_{T}g^{\prime}_{T}{(gg^{\prime})_{T}}^{-1}.$
Since $g_{T}=s(\pi(g)),g^{\prime}_{T}=s(\pi(g^{\prime})),(gg^{\prime})_{T}=s(\pi(g)\pi(g^{\prime}))$, Lemma 2.3 shows that $g_{T}g^{\prime}_{T}{(gg^{\prime})_{T}}^{-1}\in\tilde{D}(s)$ where $\tilde{D}(s)$ is the dual defect set of $s$. Hence, we have shown that $D(r)\subset\mathcal{C}(H,T)\tilde{D}(s)$ which shows that $r$ is a quasi-retraction.
[…]
It remains to verify that $s$ is a quasi-homomorphism. Note that for any $g\in G$, $r(g^{-1})\sim_{D^{-1}}r(g)^{-1}$ and for any $x,y\in G/H$, $r(s(x)),r(s(y)),r(s(xy))\in r(T)\subset D$. Since $s$ is a section, $s(y)^{-1}s(x)^{-1}s(xy)\in H$. Therefore,
$s(y)^{-1}s(x)^{-1}s(xy)$ $=r(s(y)^{-1}s(x)^{-1}s(xy))\sim_{D^{2}}r(s(y)^{-1})r(s(x)^{-1})r(s(xy))$
$\sim_{D}r(s(y)^{-1})r(s(x)^{-1})\sim_{D^{-1}}r(s(y)^{-1})r(s(x))^{-1}\sim_{D^{-1}}r(s(y)^{-1})\sim_{D^{-2}}1.$
Since $x,y\in G/H$ are arbitrary, the above formula shows that the defect of $s$ is finite, which means that $s$ is a quasi-homomorphism.
[…]
$\phi(g)=(r(g),\pi(g)),\quad\phi^{\prime}(h,x)=h\cdot s(x).$
Since $r$ is a quasi-retraction and $\pi$ is a homomorphism, it is clear that $\phi$ is a quasi-homomorphism. As for $\phi^{\prime}$, we compute directly that
$\phi^{\prime}(h_{2},x_{2})^{-1}\phi^{\prime}(h_{1},x_{1})^{-1}\phi^{\prime}(h_{1}h_{2},x_{1}x_{2})=s(x_{2})^{-1}h_{2}^{-1}s(x_{1})^{-1}h_{2}s(x_{1}x_{2})$
$=s(x_{2})^{-1}[h_{2}^{-1},s(x_{1})^{-1}]s(x_{2})s(x_{2})^{-1}s(x_{1})^{-1}s(x_{1}x_{2})\sim_{D(s)}s(x_{2})^{-1}[h_{2}^{-1},s(x_{1})^{-1}]s(x_{2})$
$=[s(x_{2})^{-1},[h_{2}^{-1},s(x_{1})^{-1}]][h_{2}^{-1},s(x_{1})^{-1}]\sim_{\mathcal{C}(H,T^{-1})}[s(x_{2})^{-1},[h_{2}^{-1},s(x_{1})^{-1}]]\sim_{\mathcal{C}(T^{-1},H)}1$
for any $(h_{1},x_{1}),(h_{2},x_{2})\in H\times G/H$. By Lemma 3.12, $\mathcal{C}(H,T^{-1})$ and $\mathcal{C}(T^{-1},H)$ are both finite since $\mathcal{C}(H,T)$ is finite. Therefore, $\phi^{\prime}$ is also a quasi-homomorphism.
It remains to show that $\phi^{\prime}=\phi^{-1}$. For any $g\in G$, $\phi^{\prime}(\phi(g))=\phi^{\prime}(r(g),\pi(g))=r(g)s(\pi(g))=g_{H}g_{T}=g$. For any $(h,x)\in H\times G/H$, $\phi(\phi^{\prime}(h,x))=(r(hs(x)),\pi(hs(x)))=(h,x)$. Thus $\phi^{\prime}=\phi^{-1}$, which completes the proof.
[…]
Remark 3.16.
In the classical Theorem 3.11, the direct product structure of $G$ guarantees that there exists a section $s:G/H\to G$ which is a homomorphism. However, different to the classical version, the finiteness of $\mathcal{C}(H,s(G/H))$ does not imply that $s$ is a quasi-homomorphism in Theorem 3.15. For example, let $G$ be the discrete Heisenberg group and $H=Z(G)$. Then $\mathcal{C}(H,s(G/H))=\{1\}$ for any normalized section $s:G/H\to G$. However, Example 3.7 (2) shows that the central extension $1\to H\to G\to G/H\to 1$ is not left quasi-split. Thus as a result of Theorem 3.15, any normalized section $s:G/H\to G$ can not be a quasi-homomorphism. By the way, Example 3.7 (1) shows that one can not drop the requirement that $\mathcal{C}(H,s(G/H))$ is finite in Theorem 3.15, either.
Corollary 3.17.
Let $H$ be a finite-index normal subgroup of $G$. Then $H$ is a quasi-retract of $G$ if and only if there exists a right coset transversal $T$ containing $1$ of $H$ in $G$ such that $\mathcal{C}(H,T)$ is finite. Moreover, if one of the above conditions holds, then $H$ is quasi-isomorphic to $G$.
Proof.
When $H$ is a finite-index normal subgroup of $G$, any section $s:G/H\to G$ is a quasi-homomorphism since $G/H$ is a finite group. Then the equivalence follows from Theorem 3.15. Moreover, if one of the above conditions holds, then Theorem 3.15 shows that $G$ is strictly quasi-isomorphic to $H\times G/H$. Since $G/H$ is a finite group, $H\times G/H$ is clearly quasi-isomorphic to $H$. Thus by Lemma 2.5, $G$ is quasi-isomorphic to $H$. ∎
In order to utilize Theorem 3.15, we need a method to transform a normal quasi-retract into a quasi-retract.
Lemma 3.18.
Let $(G,H)$ be a group pair. If there exists a finite-index subgroup $H_{0}\leq H$ such that $H_{0}$ is a normal quasi-retract of $G$, then $H$ is a quasi-retract of $G$.
Proof.
By Theorem 3.15, $H_{0}$ is a normal quasi-retract of $G$ if and only if there exists a normalized section $s:G\to H_{0}$ such that $s$ is a quasi-homomorphism and $\mathcal{C}(H_{0},s(G/H_{0}))$ is finite. Denote $T=s(G/H_{0})$ and $K=s(H/H_{0})$. Since $[H:H_{0}]<\infty$, $K\subset T$ is finite. For any $g\in G$, there is a unique decomposition $g=g_{H_{0}}g_{T}$ such that $g_{H_{0}}\in H_{0},g_{T}\in T$. Moreover, $g\in H$ if and only if $g_{T}\in K$.
Denote by $r_{0}:G\to H_{0}$ the quasi-retraction. Define a map $r:G\to H$ as follows:
$\forall g=g_{H_{0}}g_{T},r(g)=\left\{\begin{array}[]{ll}g_{H_{0}}g_{T},&\hbox{$g_{T}\in K$;}\\ g_{H_{0}},&\hbox{$g_{T}\in T\setminus K$.}\end{array}\right.$
Clearly, $r|_{H}=Id_{H}$. It remains to verify that $r$ is a quasi-homomorphism. For any two elements $g,g^{\prime}\in G$, we denote $g=g_{H_{0}}g_{T},g^{\prime}=g^{\prime}_{H_{0}}g^{\prime}_{T}$ where $g_{H_{0}},g^{\prime}_{H_{0}}\in H_{0}$ and $g_{T},g^{\prime}_{T}\in T$. By Lemma 2.4, one has that $(g_{T}g^{\prime}_{T})_{H_{0}}\in\tilde{D}(s)$ and then
(2) $(gg^{\prime})_{H_{0}}=g_{H_{0}}(g_{T}g^{\prime}_{H_{0}}g_{T}^{-1})(g_{T}g^{\prime}_{T})_{H_{0}}\sim_{\tilde{D}(s)}g_{H_{0}}g^{\prime}_{H_{0}}[{g^{\prime}_{H_{0}}}^{-1},g_{T}]\sim_{\mathcal{C}(H_{0},T)}g_{H_{0}}g^{\prime}_{H_{0}}.$
Denote $A=\mathcal{C}(H_{0},T)\tilde{D}(s)$ and $B=\mathcal{C}(H_{0},T^{-1})K^{-1}$. Then $A,B$ are both finite by Lemmas 2.3 and 3.12. The rest of the proof is completed by considering which sets $g_{T}$ and $g^{\prime}_{T}$ fall into, respectively.
Case I: $g_{T},g^{\prime}_{T}\in K\Leftrightarrow g,g^{\prime}\in H$. In this case, $gg^{\prime}\in H$. Thus $r(gg^{\prime})=gg^{\prime}=r(g)r(g^{\prime})$.
Case II: $g_{T}\in K,g^{\prime}_{T}\in T\setminus K\Leftrightarrow g\in H,g^{\prime}\notin H$. In this case, $gg^{\prime}\notin H$ and thus $(gg^{\prime})_{T}\in T\setminus K$. By the above formula (2),
$r(gg^{\prime})=(gg^{\prime})_{H_{0}}\sim_{A}g_{H_{0}}g^{\prime}_{H_{0}}=g_{H_{0}}g_{T}g^{\prime}_{H_{0}}\cdot[{g^{\prime}_{H_{0}}}^{-1},g_{T}^{-1}]g_{T}^{-1}\sim_{B}gg^{\prime}_{H_{0}}=r(g)r(g^{\prime}).$
Case III: $g_{T}\in T\setminus K,g^{\prime}_{T}\in K\Leftrightarrow g\notin H,g^{\prime}\in H$. In this case, $gg^{\prime}\notin H$ and thus $(gg^{\prime})_{T}\in T\setminus K$. By the above formula (2),
$r(gg^{\prime})=(gg^{\prime})_{H_{0}}\sim_{A}g_{H_{0}}g^{\prime}_{H_{0}}\sim_{K^{-1}}gg^{\prime}_{H_{0}}g^{\prime}_{T}=r(g)r(g^{\prime}).$
Case IV: $g_{T},g^{\prime}_{T}\in T\setminus K\Leftrightarrow g,g^{\prime}\notin H$. If $(gg^{\prime})_{T}\in K$, then $gg^{\prime}\in H$ and the above formula (2) shows that
$r(gg^{\prime})=(gg^{\prime})_{H_{0}}(gg^{\prime})_{T}\sim_{K^{-1}}(gg^{\prime})_{H_{0}}\sim_{A}g_{H_{0}}g^{\prime}_{H_{0}}=r(g)r(g^{\prime}).$
If $(gg^{\prime})_{T}\in T\setminus K$, then $gg^{\prime}\notin H$ and the above formula (2) shows that
$r(gg^{\prime})=(gg^{\prime})_{H_{0}}\sim_{A}g_{H_{0}}g^{\prime}_{H_{0}}=r(g)r(g^{\prime}).$
In summary, we conclude that $r$ is a quasi-homomorphism and the conclusion follows. ∎
Proof of Theorem 1.1.
Let $H$ be a finite-index subgroup of $G$ and $\mathrm{Core}_{G}(H)$ be the core of $H$ in $G$. Clearly $\mathrm{Core}_{G}(H)$ is a finite-index normal subgroup of $G$. Then the equivalence follows from Corollary 3.17 and the “Moreover” part follows from Lemma 3.18. ∎
Remark 3.19.
We remark that the inverse of Theorem 1.1 is not true, either. That is, if $H\leq G$ is a quasi-retract of finite index, then $\mathrm{Core}_{G}(H)$ may not be a quasi-retract of $G$. For example, let
$G=\langle a,b,c\mid a^{3}=1,ba=a^{2}b,ca=ac\rangle=\langle a\rangle\rtimes\langle b,c\rangle\cong\mathbb{Z}_{3}\rtimes F_{2}$
and $H=\langle b,c\rangle\cong F_{2}$. Clearly $H\leq G$ is a subgroup of index 3. It is easy to check that the map $r:G\to H,a\mapsto 1,b\mapsto b,c\mapsto c$ gives a retraction from $G$ to $H$. Thus $H$ is a finite-index retract of $G$.
Observe that $\mathrm{Core}_{G}(H)=H\cap aHa^{-1}\cap a^{-1}Ha$ is a proper subgroup of $H$ since $b\notin aHa^{-1}$. Hence, $[H:\mathrm{Core}_{G}(H)]\geq 2$. Suppose to the contrary that $\mathrm{Core}_{G}(H)$ is a quasi-retract of $G$. Then by Corollary 3.17, there exists a right coset transversal $T$ containing $1$ of $\mathrm{Core}_{G}(H)$ in $G$ such that $\mathrm{Core}_{G}(H)$ almost commutes with $T$. Since $[H:\mathrm{Core}_{G}(H)]\geq 2$, $T\cap H$ contains more than one element. Let $g\in T\cap H$ such that $g\notin\mathrm{Core}_{G}(H)$. Since $H\cong F_{2}$, the set $\mathcal{C}(\mathrm{Core}_{G}(H),g)$ is clearly an infinite subset of $H$. However, this is a contradiction to the finiteness of $\mathcal{C}(\mathrm{Core}_{G}(H),T)$.
In the end of this subsection, we prove Proposition 1.3 which characterize the kind of groups that each of its subgroups is a quasi-retract.
Recall that a group is non-elementary if it is not virtually abelian. An elementary group $G$ is FZ if the center of $G$ is of finite index in $G$.
Lemma 3.20.
[9, §2] A finitely generated group $G$ has finite $\mathcal{C}(G,G)$ if and only if it is FZ.
Proposition 3.21.
A finitely generated elementary group $G$ is FZ if and only if every subgroup of $G$ is a quasi-retract.
Proof.
````absent`$\Rightarrow$` ` ⇒”: Let $Z=Z(G)$ and $H$ be an arbitrary subgroup of $G$. By assumption, $[G:Z]<\infty$. Since $Z$ is an abelian group, Lemma 3.6 (5) shows that $H\cap Z$ is a quasi-retract of $Z$. Since $Z$ commutes with $G$, Corollary 3.17 shows that $Z$ is a quasi-retract of $G$. Thus, as a composition, $H\cap Z$ is a normal quasi-retract of $G$. Note that $H/H\cap Z\cong HZ/Z$. This implies that $[H:H\cap Z]<\infty$. By Lemma 3.18, one gets that $H$ is a quasi-retract of $G$.
````absent`$\Leftarrow$` ` ⇐”: Since $G$ is virtually abelian, we can choose a finite-index, normal and abelian subgroup $A$ of $G$. By assumption, $A$ is a quasi-retract of $G$. By Corollary 3.17, there exists a finite coset transversal $T$ containing $1$ of $A$ in $G$ such that $\mathcal{C}(A,T)$ is finite. Since $T$ is a coset transversal of $A$, any element $g$ has a unique decomposition $g=g_{A}g_{T}$ where $g_{A}\in A,g_{T}\in T$. For any $g=g_{A}g_{T},g^{\prime}=g^{\prime}_{A}g^{\prime}_{T}\in G$, we have
$gg^{\prime}=g_{A}g_{T}g^{\prime}_{A}g^{\prime}_{T}\sim_{T}g_{A}g_{T}g^{\prime}_{A}=g_{A}g^{\prime}_{A}[{g^{\prime}_{A}}^{-1},g_{T}]g_{T}\sim_{\mathcal{C}(A,T)T}g_{A}g^{\prime}_{A}$
and similarly
$g^{\prime}g=g^{\prime}_{A}g^{\prime}_{T}g_{A}g_{T}\sim_{T}g^{\prime}_{A}g^{\prime}_{T}g_{A}=g^{\prime}_{A}g_{A}[{g_{A}}^{-1},g^{\prime}_{T}]g^{\prime}_{T}\sim_{\mathcal{C}(A,T)T}g^{\prime}_{A}g_{A}=g_{A}g^{\prime}_{A}.$
By combining both two formulas, we get that $\mathcal{C}(G,G)=\{[g,g^{\prime}]:g,g^{\prime}\in G\}$ is finite. Then the conclusion follows from Lemma 3.20. ∎