### v3

# Quasi-retracts of groups

Abstract.
In this paper, we study a special class of quasi-homomorphisms, i.e. quasi-retractions from a group to its subgroups. We first give some algebraic and geometric properties of quasi-retracts and then propose a theory of quasi-split short exact sequences of groups. Later, we establish a connection between quasi-homomorphisms and induced quasi-actions. Finally, we give some geometric applications of quasi-homomorphisms, including normal quasi-retracts inherit cobounded actions on hyperbolic spaces, properties (QFA), (QT’) and (PH’) are all stable under left quasi-split group extensions, quasi-isomorphic groups have isomorphic hyperbolic structures, and so on.
Key words and phrases:
Quasi-homomorphisms, quasi-retracts, quasi-isomorphisms, quasi-actions, hyperbolic structures
1. Introduction
1.1. Quasi-homomorphisms and quasi-retracts
Let $G$ be a group and $H$ be a group equipped with a proper left-invariant metric $d$ (e.g., a finitely generated group, equipped with a word metric). A map $\phi:G\to H$ is called a quasi-homomorphism if there exists a constant $C>0$ such that
$d(\phi(xy),\phi(x)\phi(y))\leq C$
for all $x,y\in G$. For two maps $\phi,\phi^{\prime}:G\to H$, we define
$\mathrm{dist}(\phi,\phi^{\prime}):=\sup_{x\in G}d(\phi(x),\phi^{\prime}(x)).$
Two quasi-homomorphisms $\phi$ and $\phi^{\prime}$ are called equivalent if $\mathrm{dist}(\phi,\phi^{\prime})<\infty$. In the case when $H$ is discrete (and in this paper we limit ourselves only to this class of groups), $\phi$ is a quasi-homomorphism if and only if the set of defects of $\phi$
$D(\phi)=\{\phi(y)^{-1}\phi(x)^{-1}\phi(xy):x,y\in G\}$
is finite. And two quasi-homomorphisms $\phi,\phi^{\prime}$ are equivalent if and only if $\{\phi(x)^{-1}\phi^{\prime}(x):x\in G\}$ is finite. A quasi-homomorphism with values in $\mathbb{Z}$ (or $\mathbb{R}$, equipped with the standard metric) is called a quasimorphism.
The concept of quasi-homomorphisms goes back to Ulam [32, Chapter 6], who asked if they are close to group homomorphisms. Later, Fujiwara-Kapovich [12] studied this concept to figure out why it is so “hard to construct” quasi-homomorphisms to noncommutative groups which are neither homomorphisms, nor come from quasi-homomorphisms with commutative targets, provided that $H$ is a discrete group. They showed that every quasi-homomorphism is constructible [12, Theorem 1.2].
In this paper, we study a special class of quasi-homomorphisms, i.e. quasi-homomorphisms from a group to its subgroups.
For convenience, we say $(G,H)$ is a group pair if $G$ is a discrete group and $H\leq G$ is a subgroup. Let $(G,H)$ be a group pair. Classically, $H$ is called a retract of $G$ if there exists a homomorphism $r:G\to H$ such that $r$ is left-inverse to the inclusion map $\iota:H\to G$. This homomorphism $r$ is called a retraction from $G$ to $H$. The study of retractions have been lasting for decades in Geometric Group Theory and relates to many other topics. For example, [28, 22, 31] studied retracts of free groups to characterize fixed subgroups of automorphisms of free groups and [23, 19] introduced the notion of virtual retractions to study the geometry and topology of hyperbolic manifolds and discrete groups.
In this paper, we define a subgroup $H$ of $G$ to be a quasi-retract [^1] This word “quasi-retraction” was first used by Alonso in [2] to denote a certain subordinate relationship between metric spaces. if there exists a quasi-homomorphism $r:G\to H$ such that $r\circ\iota$ is equivalent to $Id_{H}$ where $\iota:H\to G$ is the inclusion map. This quasi-homomorphism $r$ is called a quasi-retraction from $G$ to $H$. By definition, retracts are quasi-retracts.
A relationship between quasimorphisms and quasi-retracts is that a group admits an unbounded quasimorphism if and only if it contains an infinite cyclic quasi-retract (cf. Lemma 3.3).
Another thing worth mentioning is that in [15], Hartnick-Schiweitzer defined a map $\phi:G\to H$ to be a quasi-homomorphism if and only if any quasimorphism on $H$ can be pull-backed to a quasimorphism on $G$ by composing $\phi$. To make a distinction, we shall call such a quasi-homomorphism an HS-quasi-homomorphism. Quasi-homomorphisms and HS-quasi-homomorphisms are not equivalent in general; see [12, Subsection 9.3] for a detailed discussion about different quasi-homomorphisms. For a group pair $(G,H)$, Hartnick-Schiweitzer also defined an HS-quasi-homomorphism $r:G\to H$ to be an HS-quasi-retraction if the composition $r\circ\iota$ is equivalent to $Id_{H}$ where $\iota:H\to G$ is the inclusion map.
1.2. Extension problem
In [17], Hull-Osin considered the following “extension problem”:
Question 1.
Under what conditions can a 1-quasi-cocycle (eg. a quasimorphism) on a subgroup be extended to the whole group?
They solved the problem by assuming the subgroup is hyperbolically embedded. Roughly speaking, hyperbolically embedded subgroups generalize the concept of parabolic subgroups in relatively hyperbolic groups; see [9] for details about hyperbolically embedded subgroups. Moreover, under the same assumption, Frigerio-Pozzetti-Sisto [10] solved a “higher-dimensional extension problem”.
From the definition, any quasimorphism on a quasi-retract can be extended via this quasi-retraction to a quasimorphism on the whole group. So quasi-retracts naturally solve the “extension problem” about quasimorphisms.
The first result of this paper is to explore the relationship between hyperbolically embedded subgroups and quasi-retracts. We denote by $\mathbb{F}_{n}$ the free group of rank $n$. A quasi-retract is called nontrivial if it is neither a finite subgroup nor the whole group.
Theorem 1.1.
  1. (1)
    A finite-by-$\mathbb{Z}^{m}$ hyperbolically embedded subgroup is always a quasi-retract.
  2. (2)
    A subgroup of $\mathbb{F}_{2}$ is a nontrivial quasi-retract if and only if it is a cyclic group. Moreover, there exists a hyperbolically embedded subgroup of $\mathbb{F}_{2}$ which is not a quasi-retract.
Remark 1.2.
  1. (1)
    It is well known and easy to prove that every infinite virtually cyclic group is either finite-by-(infinite cyclic) or finite-by-(infinite dihedral). Hence, for a group without involutions, any virtually cyclic subgroup is finite-by-cyclic.
  2. (2)
    There are several typical examples of groups containing virtually cyclic hyperbolically embedded subgroups (cf. [17]):
    • $G$ is the mapping class group of a punctured closed orientable surface and $H\leq G$ is the maximal virtually cyclic subgroup containing a pseudo-Anosov element;
    • $G=\mathrm{Out}(\mathbb{F}_{n})$ and $H\leq G$ is the maximal virtually cyclic subgroup containing a fully irreducible automorphism;
    • $G$ is a group acting properly on a proper CAT(0) space and $H\leq G$ is the maximal virtually cyclic subgroup containing a rank-1 element.
1.3. Quasi-split short exact sequence
Let
(*) $1\to H\xrightarrow{\iota}G\xrightarrow{\pi}Q\to 1$
be a short exact sequence of groups. Classically, the short exact sequence (*1.3) is left split if there exists a retraction from $G$ to $H$. It is right split if there exists a section $s:Q\to G$ such that $s$ is a homomorphism. Similarly, we say the short exact sequence (*1.3) is left quasi-split if there exists a quasi-retraction from $G$ to $H$. It is right quasi-split if there exists a section $s:Q\to G$ such that $s$ is a quasi-homomorphism. We remark that the notion of right quasi-split here is the same as the notion of quasisplit in [12, Subsection 2.4.1].
Theorem 1.3 (Classical version).
Let $1\to H\xrightarrow{\iota}G\xrightarrow{\pi}Q\to 1$ be a short exact sequence of groups. The following are equivalent:
  1. (1)
    The short exact sequence is left split. Or equivalently, $H$ is a retract of $G$.
  2. (2)
    There exists a section $s:Q\to G$ such that $s$ is a homomorphism and $H$ commutes with $s(Q)$.
  3. (3)
    There is an isomorphism $\phi:G\to H\times Q$ such that the diagram commutes:
[…]
    where the bottom row is the short exact sequence for a direct product.
A section $s$ is normalized if $s(1)=1$. For two subsets $K,T$ of a group $G$, we define $\mathcal{C}(K,T)=\{[k,t]=ktk^{-1}t^{-1}:k\in K,t\in T\}$. In this case, $K$ and $T$ are commutative (resp. almost commutative) if and only if $\mathcal{C}(K,T)=\{1\}$ (resp. $\mathcal{C}(K,T)$ is finite).
For two discrete groups $G,G^{\prime}$, a quasi-homomorphism $\phi:G\to G^{\prime}$ is a quasi-isomorphism if it admits a quasi-inverse, i.e., a quasi-homomorphism $\phi^{\prime}:G^{\prime}\to G$ such that $\mathrm{dist}(\phi^{\prime}\circ\phi,Id_{G})<\infty$ and $\mathrm{dist}(\phi\circ\phi^{\prime},Id_{G^{\prime}})<\infty$. A quasi-isomorphism is strict if $\phi^{\prime}=\phi^{-1}$. The second result of this paper is an analogue of Theorem 1.3.
Theorem 1.4 (Theorem 4.5).
Let $1\to H\xrightarrow{\iota}G\xrightarrow{\pi}Q\to 1$ be a short exact sequence of groups. The following are equivalent:
  1. (1)
    The short exact sequence is left quasi-split. Or equivalently, $H$ is a quasi-retract of $G$.
  2. (2)
    There exists a normalized section $s:Q\to G$ such that $s$ is a quasi-homomorphism and $H$ almost commutes with $s(Q)$.
  3. (3)
    There is a strict quasi-isomorphism $\phi:G\to H\times Q$ such that the diagram commutes:
[…]
    where the bottom row is the short exact sequence for a direct product.
Remark 1.5.
The direction $(2)\Rightarrow(3)$” in Theorem 1.4 is also an analogue of [2, Proposition A.2] which states that by adding some finiteness conditions to the group extension (*1.3), $G$ is quasi-isometric to $H\times Q$. We will see later that quasi-isomorphisms are quasi-isometries for finitely generated groups.
Central extensions
An important class of short exact sequences comes from central extensions. Let
(**) $1\to Z\xrightarrow{\iota}G\xrightarrow{\pi}Q\to 1$
be a central extension of groups, where $Q$ and $Z$ (hence, $G$) are finitely generated. Any such extension defines a cohomology class $\omega\in H^{2}(Q,Z)$ which will be called the Euler class of the extension. It is well known that the Euler class completely determines the isomorphism class of a central extension, and it is natural to investigate which geometric features it encodes.
We say that a class $\omega\in H^{2}(Q,Z)$ is bounded if it lies in the image of the comparison map $H^{2}_{b}(Q,Z)\to H^{2}(Q,Z)$, i.e. if it can be described by a bounded cocycle. In [12], Fujiwara-Kapovich proved that the central extension (**1.3) has a bounded Euler class if and only if it is right quasi-split. By Theorem 1.4, this is also equivalent to saying that $Z$ is a quasi-retract of $G$. For more equivalent conditions for (**1.3) to have a bounded Euler class, we refer the readers to [30, Lemma 5.6].
Following [13], we say that the extension of finitely generated groups
$1\to Z\xrightarrow{\iota}G\xrightarrow{\pi}Q\to 1$
is quasi-isometrically trivial if there exists a quasi-isometry $\phi:G\to Z\times Q$ such that the following diagram commutes, up to bounded error:
[…]
Here $\pi_{2}:Z\times Q\to Q$ is the projection on the second factor, and, for clarity, the diagram commuting up to bounded error means that given a word metric $d_{Q}$ on $Q$ we have $\sup_{g\in G}d_{Q}(\pi(g),\pi_{2}(\phi(g)))<+\infty$.
Gersten proved in [13, §3] that if the Euler class of the central extension is bounded, then the extension is quasi-isometrically trivial. As for the converse direction, Frigerio-Sisto investigated in [11] when a quasi-isometrically trivial central extension has a bounded Euler class. They proved that a central extension is quasi-isometrically trivial if and only if it has a weakly bounded Euler class. As an analogue, Theorem 1.4 implies that
Corollary 1.6.
The central extension (**1.3) has a bounded Euler class if and only if it is it is “strictly quasi-isomorphically trivial”, i.e. there is a strict quasi-isomorphism $\phi:G\to Z\times Q$ such that the diagram commutes:
[…]
where the bottom row is the short exact sequence for a direct product.
Since bounded Euler classes imply weakly bounded Euler classes, we will see in Lemma 2.10 that quasi-isomorphisms imply quasi-isometries for finitely generated groups.
The following example presents an important class of normal quasi-retracts. For the definition of amalgamated direct products, we refer the readers to Definition 3.12.
Example 1.7.
Let $\Sigma$ be a closed surface with finitely many marked points. Let $C$ be a separating essential simple closed curves in $\Sigma$. Let $\mathrm{PMCG}(\Sigma)$ be the pure mapping class group of $\Sigma$ and $\mathrm{PMCG}(\Sigma,C)$ denote the stabilizer in $\mathrm{PMCG}(\Sigma)$ of $C$. Since $C$ separates $\Sigma$, $C$ divides $\Sigma$ into two subsurfaces $\Sigma_{-},\Sigma_{+}$ with a common boundary component $C$. It is well-known and easy to prove that $\mathrm{PMCG}(\Sigma,C)$ is an amalgamated direct product of $\mathrm{PMCG}(\Sigma_{-})$ and $\mathrm{PMCG}(\Sigma_{+})$. Suppose that $\Sigma_{+}^{\prime}$ is the surface obtained from $\Sigma_{+}$ by capping the boundary component $C$ with a once-marked disk. By [12, Proposition 6.4], the following central extension
$1\to\langle T_{C}\rangle\to\mathrm{PMCG}(\Sigma_{+})\xrightarrow{Cap}\mathrm{PMCG}(\Sigma_{+}^{\prime})\to 1$
has a bounded Euler class where $T_{C}$ is the Dehn twist around $C$. As a result of Lemma 3.14 (3), $\mathrm{PMCG}(\Sigma_{-})$ is a normal quasi-retract of $\mathrm{PMCG}(\Sigma,C)$.
1.4. Induced quasi-actions
In [21], Manning defined group quasi-actions as a natural coarse generalization of isometric group actions. Specifically, a $(\lambda,\epsilon)$-quasi-action of a group $G$ on a metric space $X$ is a map $\rho:G\times X\to X$, denoted $\rho(g,x)\mapsto gx$, so that the following hold:
  • (i)
    for each $g$, $\rho(g,-):X\to X$ is a $(\lambda,\epsilon)$-quasi-isometry;
  • (ii)
    for each $x\in X$ and $g,h\in G$, we have
    $d(g(hx),(gh)x)=d(\rho(g,\rho(h,x)),\rho(gh,x))\leq\epsilon.$
(Note that $\lambda$ and $\epsilon$ must be independent of $g$ and $h$.) We call a quasi-action cobounded if, for every $x\in X$, the map $\rho(-,x):G\to X$ is $\epsilon^{\prime}$-coarsely surjective for some $\epsilon^{\prime}\geq 0$.
An important geometric application of quasi-homomorphisms is that they can induce new quasi-actions on metric spaces. A key lemma is as follows.
Lemma 1.8 (Lemma 5.15).
Let $\phi:G\to H$ be a coarsely surjective quasi-homomorphism. Suppose $H$ admits a cobounded quasi-action on a metric space $X$. Then the map $\rho:G\times X\to X$ defined by $\rho(g,x):=\phi(g)x$ is an induced quasi-action of $G$ on $X$ by $\phi$.
Moreover, suppose $H$ admits another cobounded quasi-action on a metric space $Y$. Then the quasi-action of $H$ on $X$ is quasi-conjugate to the quasi-action of $H$ on $Y$ if and only if the induced quasi-action of $G$ on $X$ by $\phi$ is quasi-conjugate to the induced quasi-action of $G$ on $Y$ by $\phi$.
Property (QFA), (QT) and (PH)
To apply Lemma 1.8, we first introduce some geometric properties of groups acting isometrically on hyperbolic spaces.
In [21], Manning defined a group $G$ to have property (QFA) if for every quasi-action of $G$ on any tree $X$, there is some $x\in X$ so that the orbit $Gx$ has finite diameter. This property is analogous to Serre’s property (FA). A typical class of groups with property (QFA) are $\mathrm{SL}_{n}(\mathbb{Z})$ for $n\geq 3$ [21].
In [5], Bestvina-Bromberg-Fujiwara defined a finitely generated group $G$ to have property (QT) if $G$ acts isometrically on a finite product of quasi-trees equipped with $\ell^{1}$-metric such that the orbit map is a quasi-isometric embedding. They also showed that residually finite hyperbolic groups and mapping class groups have property (QT). Following [29], a finitely generated group $G$ has property (QT’) if it has property (QT) and the action on a finite product of quasi-trees is induced by a diagonal action. By a result of Button [7], a finitely generated group $G$ with property (QT) virtually has property (QT’), i.e. $G$ contains a finite-index subgroup $G_{0}$ which has property (QT’).
In [30], Tao and the author defined a group $G$ to have property (PH’) if there exist finitely many hyperbolic spaces $X_{1},\ldots,X_{n}$ on which $G$ acts coboundedly and the induced diagonal action of $G$ on the product $\prod_{i=1}^{n}X_{i}$ with $\ell^{1}$-metric is proper. A group $G$ has property (PH) if $G$ virtually has property (PH’). Some typical examples of groups with property (PH) include finitely generated abelian groups, Coxeter groups, virtually colorable hierarchically hyperbolic groups and most 3-manifold groups. As shown in [30, Lemma 2.2], if a finitely generated group has property (QT) or (QT’), then it has property (PH) or (PH’), respectively. We refer the readers to [30, Figure 1] for an illustration of relations between these properties.
By combining Lemma 1.8 and a result of Manning (cf. Proposition 5.20), we can obtain some quasi-isomorphism invariants for discrete groups. Since some of them are not quasi-isometry invariants, we know that quasi-isomorphisms are strictly stronger equivalence relations than quasi-isometries for finitely generated groups.
Proposition 1.9.
  1. (1)
    Almost commutativity is a quasi-isomorphism invariant for discrete groups.
  2. (2)
    Property (QFA) is a quasi-isomorphism invariant for finitely generated groups.
  3. (3)
    Property (PH’) is a quasi-isomorphism invariant for discrete groups.
  4. (4)
    Property (QT’) is a quasi-isomorphism invariant for finitely generated groups.
In [30], Tao and the author proved the stability of properties (QT), (QT’), (PH) and (PH’) under left quasi-split central extensions. Here, we give a partial generalization of [30, Theorems 1.7, 1.8]. A group is called FZ if its center is of finite index.
Theorem 1.10.
Let $1\to H\to G\to Q\to 1$ be a short exact sequence of finitely generated groups which is left quasi-split. Then
  1. (1)
    $G$ has property (QFA) if and only if both $H$ and $Q$ have property (QFA).
  2. (2)
    $G$ has property (PH’) or (QT’) if and only if both $H$ and $Q$ have property (PH’) or (QT’), respectively.
  3. (3)
    suppose that $H$ is FZ. Then $G$ has property (PH) or (QT) if and only if $Q$ has property (PH) or (QT), respectively.
Hyperbolic structures
For every group $G$, Abbott-Balasubramanya-Osin [1] defined the set of hyperbolic structures on $G$, denoted by $\mathcal{H}(G)$, which consists of equivalence classes of (usually infinite) generating sets of $G$ such that the corresponding Cayley graph is hyperbolic; two generating sets of $G$ are equivalent if the corresponding word metrics on $G$ are bi-Lipschitz equivalent. See §5.1 or [1] for more details about hyperbolic structures on groups.
Finally, we obtain a relationship on hyperbolic structures between a group and its quasi-retracts.
Proposition 1.11.
  1. (1)
    Let $\phi:G\to H$ be a coarsely surjective quasi-homomorphism. Then $\mathcal{H}(H)$ embeds in $\mathcal{H}(G)$ as a sub-poset.
  2. (2)
    Let $G$ and $G^{\prime}$ be two discrete quasi-isomorphic groups. Then $\mathcal{H}(G)\cong\mathcal{H}(G^{\prime})$.
Structure of the paper
The paper is organized as follows. Section 2 is devoted to defining some notations and studying some elementary properties of quasi-homomorphisms and quasi-isomorphisms. In Section 3, we first study some basic properties of quasi-retractions and then characterize the quasi-retracts of torsion-free hyperbolic groups and some more general groups. In Section 4, we introduce the classical theory of split short exact sequences and prove Theorem 1.4. By utilizing Theorem 1.4, we complete the proof of Theorem 1.1. In Section 5, we recall some basic material about group actions on hyperbolic spaces, and then prove Proposition 1.9, Theorem 1.10 and Proposition 1.11 in turn.
Acknowledgments
We are grateful to Zhenguo Huangfu and Bingxue Tao for many helpful suggestions on the first draft. R. W. is supported by NSFC No.12471065 & 12326601 and in part by Science and Technology Commission of Shanghai Municipality (No. 22DZ2229014).
2. Preliminaries
2.1. Definitions and notations
Let $H$ be a group with a proper left-invariant metric $d$. For a subset $A\subset H$ and a constant $C>0$, we denote $\mathcal{N}_{C}(A):=\{h\in H:d(h,A)\leq C\}$.
Let $S\subset H$ be a finite subset. For convenience, we will always write
$h\sim_{S}h^{\prime}$
if $h=h^{\prime}s$ with $h,h^{\prime}\in H,s\in S$. For example, for a quasi-homomorphism $\phi:G\to H$ with a defect set $D=D(\phi)$, by the definition,
[…]
for $a,b,c\in G$.
For a subset $D$ of a group $H$ and $n\geq 2$ we will use the notation $D^{n}$ to denote the subset of $H$ consisting of products of at most $n$ elements of $D$. More generally, for two subsets $A,B\subset H$ we let
$A\cdot B=\{ab:a\in A,b\in B\}.$
We will use the notation $D^{-1}$ for the set of inverses of elements of $D$. Then
$h\sim_{D}h^{\prime}\Longleftrightarrow h^{\prime}\sim_{D^{-1}}h.$
Given a group $H$ and its subgroup $A$ we let $N_{H}(A)$ and $Z_{H}(A)$ denote the normalizer and the centralizer of $A$ in $H$ respectively.
Given a group $G$ with a finite generating set $S$, we denote by $d_{S}$ the word metric on $G$.
Given a quasi-homomorphism $\phi:G\to H$, we denote by $\Delta_{\phi}$ the defect subgroup of $H$ which is finitely generated by $D(\phi)$.
2.2. Elementary properties of quasi-homomorphisms
We collect some basic facts about quasi-homomorphisms. See [12, Subsections 2.2, 2.3, 3.1] for their proofs.
Lemma 2.1.
Let $\phi:G\to H$ be a quasi-homomorphism with $D=D(\phi)$.
  1. (1)
    For any $x_{1},\cdots,x_{n}\in G$, $\phi(x_{1}\cdots x_{n})\sim_{D^{n-1}}\phi(x_{1})\cdots\phi(x_{n})$.
  2. (2)
    For any $x\in G$, $\phi(x)^{-1}\sim_{D^{2}}\phi(x^{-1})$.
  3. (3)
    For any $h\in\phi(G)$, $h^{-1}Dh\subset D^{2}D^{-1}$.
  4. (4)
    $\phi(G)\cup\phi(G)^{-1}\subset N_{H}(\Delta_{\phi})$ and the induced map $\bar{\phi}:G\to N_{H}(\Delta_{\phi})/\Delta_{\phi}$ is a homomorphism.
  5. (5)
    There exists a constant $C$ such that $\phi(G)\subset\mathcal{N}_{C}(Z_{H}(\Delta_{\phi}))$.
As a corollary, we have
Lemma 2.2.
Let $\phi:G\to H$ be a surjective quasi-homomorphism. Then the defect subgroup $\Delta_{\phi}$ is a virtually abelian normal subgroup of $H$.
Proof.
Since $\phi$ is surjective, Lemma 2.1 (4) shows that $H=\phi(G)=N_{H}(\Delta_{\phi})$. Thus $\Delta_{\phi}\lhd H$. By Lemma 2.1 (5), there exists a constant $C$ such that $\phi(G)=H\subset\mathcal{N}_{C}(Z_{H}(\Delta_{\phi}))$. This implies that $H:Z_{H}(\Delta_{\phi})]<\infty$. Denote $\Delta_{\phi}^{\prime}=\Delta_{\phi}\cap Z_{H}(\Delta_{\phi})$. Note that $\Delta_{\phi}^{\prime}$ is an abelian group. According to the isomorphism theorem of groups, $\Delta_{\phi}:\Delta_{\phi}^{\prime}]\leq[H:Z_{H}(\Delta_{\phi})]<\infty$, which implies that $\Delta_{\phi}$ is virtually abelian. ∎
Lemma 2.3.
Let $\phi:G\to H$ be a quasi-homomorphism with $D=D(\phi)$. For any $x,y\in G$, $\phi(x^{-1}y)\sim_{D}\phi(x^{-1})\phi(y)\sim_{(D^{2}D^{-1})^{-2}}\phi(x)^{-1}\phi(y)$.
Proof.
By Lemma 2.1 (2), there exist $a_{1},a_{2}\in D^{-1}$ such that $\phi(x^{-1})=\phi(x)^{-1}a_{1}a_{2}$. Then by Lemma 2.1 (3), one has
$\phi(x^{-1})\phi(y)=\phi(x)^{-1}a_{1}a_{2}\phi(y)=\phi(x)^{-1}\phi(y)\cdot\phi(y)^{-1}a_{1}\phi(y)\cdot\phi(y)^{-1}a_{2}\phi(y)\sim_{(D^{2}D^{-1})^{-2}}\phi(x)^{-1}\phi(y).$
Recall that the commutator set of a group $G$ is defined as: $\mathcal{C}(G,G)=\{[x,y]:x,y\in G\}$.
Lemma 2.4.
Let $\phi:G\to H$ be a quasi-homomorphism with $D=D(\phi)$. Then there exists a constant $C$ such that $\phi(\mathcal{C}(G,G))\subset\mathcal{N}_{C}(\mathcal{C}(H,H))$.
Proof.
For any $x,y\in G$, Lemma 2.3 and Lemma 2.1 (1), (2) show that
$\phi(xyx^{-1}y^{-1})$ $\sim_{D^{3}}\phi(x)\phi(y)\phi(x^{-1})\phi(y^{-1})$
$\sim_{(D^{2}D^{-1})^{-2}}\phi(x)\phi(y)\phi(x)^{-1}\phi(y^{-1})$
$\sim_{D^{-2}}\phi(x)\phi(y)\phi(x)^{-1}\phi(y)^{-1}\in\mathcal{C}(H,H).$
Then the conclusion follows since $D$ is a finite set in $H$. ∎
For a map $\phi:G\to H$, one can define another kind of defect set of $\phi$ as
$\tilde{D}(\phi)=\{\phi(x)\phi(y)\phi(xy)^{-1}:x,y\in G\}.$
Lemma 2.5.
[16, Proposition 2.3] A map $\phi:G\to H$ is a quasi-homomorphism if and only if $\tilde{D}(\phi)$ is finite.
In general, a bounded perturbation of a quasi-homomorphism is not necessarily a quasi-homomorphism. However, if the perturbation is achieved through finitely many central elements, then it remains a quasi-homomorphism.
Lemma 2.6.
Let $\phi:G\to H$ be a quasi-homomorphism and $\phi^{\prime}:G\to H$ be a map. If there exists a finite subset $A\subset Z(H)$ such that $\phi^{\prime}(g)\sim_{A}\phi(g)$ for any $g\in G$, then $\phi^{\prime}$ is a quasi-homomorphism equivalent to $\phi$.
Proof.
Let $D=D(\phi)$ be the defect set of $\phi$. It is easy to see that $D(\phi^{\prime})\subset A^{-2}DA$ and thus $\phi^{\prime}$ is also a quasi-homomorphism. ∎
Let $(G,H)$ be a group pair and $T$ be a right coset transversal of $H$ in $G$. It is clear that any element $g\in G$ can be written uniquely as $g=g_{H}g_{T}$ where $g_{H}\in H$ and $g_{T}\in T$. The following simple lemma will be used in later proofs.
Lemma 2.7.
Let $(G,H)$ be a group pair and $T$ be a right coset transversal of $H$ in $G$. If $H$ is a normal subgroup of $G$, then the product of any two elements $g,g^{\prime}\in G$ satisfies that $(gg^{\prime})_{H}=g_{H}(g_{T}g^{\prime}_{H}g_{T}^{-1})(g_{T}g^{\prime}_{T})_{H}$ and $(gg^{\prime})_{T}=(g_{T}g^{\prime}_{T})_{T}$. Moreover, $(g_{T}g^{\prime}_{T})_{H}\in\tilde{D}(s)$ where $s:G/H\to G$ is a section such that $T=s(G/H)$.
Proof.
It is straightforward to calculate that
$gg^{\prime}=g_{H}g_{T}g^{\prime}_{H}g^{\prime}_{T}=g_{H}(g_{T}g^{\prime}_{H}g_{T}^{-1})g_{T}g^{\prime}_{T}=g_{H}(g_{T}g^{\prime}_{H}g_{T}^{-1})(g_{T}g^{\prime}_{T})_{H}(g_{T}g^{\prime}_{T})_{T}.$
Since $H\lhd G$, $g_{T}g^{\prime}_{H}g_{T}^{-1}\in H$. By the uniqueness of decomposition, one gets that $(gg^{\prime})_{H}=g_{H}(g_{T}g^{\prime}_{H}g_{T}^{-1})(g_{T}g^{\prime}_{T})_{H}$ and $(gg^{\prime})_{T}=(g_{T}g^{\prime}_{T})_{T}$. Moreover, let $\pi:G\to G/H$ be the quotient map and $s:G/H\to G$ be a section such that $T=s(G/H)$. This means that for any $g=g_{H}g_{T}\in G$, $s(\pi(g))=s(\pi(g_{T}))=g_{T}$. Then $s(\pi(gg^{\prime}))=(gg^{\prime})_{T}=(g_{T}g^{\prime}_{T})_{T}$. Finally, it follows from $g_{T}g^{\prime}_{T}=(g_{T}g^{\prime}_{T})_{H}(g_{T}g^{\prime}_{T})_{T}$ that
$(g_{T}g^{\prime}_{T})_{H}=g_{T}g^{\prime}_{T}(g_{T}g^{\prime}_{T})_{T}^{-1}=s(\pi(g))s(\pi(g^{\prime}))s(\pi(gg^{\prime}))^{-1}\in\tilde{D}(s).$
2.3. Elementary properties of quasi-isomorphisms
Recall that a quasi-homomorphism $\phi:G\to H$ is a quasi-isomorphism if there exists a quasi-homomorphism $\phi^{\prime}:H\to G$, called the quasi-inverse of $\phi$ such that $\mathrm{dist}(\phi^{\prime}\circ\phi,Id_{G})<\infty$ and $\mathrm{dist}(\phi\circ\phi^{\prime},Id_{H})<\infty$. Moreover, a quasi-isomorphism is called strict if $\phi^{\prime}=\phi^{-1}$.
Lemma 2.8.
Quasi-isomorphism and strict quasi-isomorphism are two equivalent relations on discrete groups.
Proof.
Both reflexivity and symmetry are easy to see. We only need to verify the transitivity. Let $\phi_{1}:G\to H$ and $\phi_{2}:H\to K$ be two quasi-isomorphisms between discrete groups. By definition, there exist two quasi-homomorphisms $\phi_{1}^{\prime}:H\to G$ and $\phi_{2}^{\prime}:K\to H$ and four finite subsets $A_{1}\subset G,B_{1},A_{2}\subset H,B_{2}\subset K$ such that for any $g\in G,h\in H,k\in K$,
$\phi_{1}^{\prime}(\phi_{1}(g))\sim_{A_{1}}g,\quad\phi_{1}(\phi_{1}^{\prime}(h))\sim_{B_{1}}h,\quad\phi_{2}^{\prime}(\phi_{2}(h))\sim_{A_{2}}h,\quad\phi_{2}(\phi_{2}^{\prime}(k))\sim_{B_{2}}k.$
It follows that
$\phi_{2}^{\prime}\circ\phi_{2}(\phi_{1}(g))\sim_{A_{2}}\phi_{1}(g),\quad\phi_{1}\circ\phi_{1}^{\prime}(\phi_{2}^{\prime}(k))\sim_{B_{1}}\phi_{2}^{\prime}(k).$
By composing $\phi_{1}^{\prime}$ or $\phi_{2}$ on the left, we get
(1) $\phi_{1}^{\prime}\circ\phi_{2}^{\prime}\circ\phi_{2}(\phi_{1}(g))\sim_{\phi_{1}^{\prime}(A_{2})D(\phi_{1}^{\prime})}\phi_{1}^{\prime}(\phi_{1}(g))\sim_{A_{1}}g$
and
(2) $\phi_{2}\circ\phi_{1}\circ\phi_{1}^{\prime}(\phi_{2}^{\prime}(k))\sim_{\phi_{2}(B_{1})D(\phi_{2})}\phi_{2}(\phi_{2}^{\prime}(k))\sim_{B_{2}}k.$
Denote $\phi=\phi_{2}\circ\phi_{1}:G\to K$ and $\phi^{\prime}=\phi_{1}^{\prime}\circ\phi_{2}^{\prime}:K\to G$ as two composite maps. They are both quasi-homomorphisms since a composition of quasi-homomorphisms is still a quasi-homomorphism. The above formulas (1) and (2) show that both $\{g^{-1}\phi^{\prime}(\phi(g)):g\in G\}$ and $\{k^{-1}\phi(\phi^{\prime}(k)):k\in K\}$ are finite. Therefore, $\phi,\phi^{\prime}$ are quasi-isomorphisms, which shows that quasi-isomorphism is an equivalent relation on discrete groups. For the other relation, the same proof works. ∎
A group $G$ is called almost commutative if $\mathcal{C}(G,G)$ is finite. As a corollary of Lemma 2.4 and Lemma 2.8, we have
Lemma 2.9.
Almost commutativity is a quasi-isomorphism invariant for discrete groups.
In the field of Geometric Group Theory, people always care about quasi-isometry invariants for finitely generated groups. The following lemma implies that quasi-isometry invariants are also quasi-isomorphism invariants.
Lemma 2.10.
For finitely generated groups equipped with word metrics, quasi-isomorphisms are quasi-isometries.
Proof.
Let $G,G^{\prime}$ be two finitely generated groups with finite generating sets $S,S^{\prime}$, respectively. Let $\phi:G\to G^{\prime}$ be a quasi-isomorphism and $\phi^{\prime}:G^{\prime}\to G$ be its quasi-inverse. By definition, there exists a constant $C>0$ such that $\sup_{g\in G}d_{S}(\phi^{\prime}(\phi(g)),g)\leq C$ and $\sup_{g^{\prime}\in G^{\prime}}d_{S^{\prime}}(\phi(\phi^{\prime}(g^{\prime})),g^{\prime})\leq C$. The second inequality shows that $G^{\prime}\subseteq\mathcal{N}_{C}(\phi(G))$. Thus it remains to show that $\phi:G\to G^{\prime}$ is a quasi-isometric embedding.
Let $D=D(\phi)$ and $D^{\prime}=D(\phi^{\prime})$ be the corresponding defect sets of $\phi$ and $\phi^{\prime}$, respectively. Denote $C_{1}=\max_{x\in D^{\prime}}d_{S}(1,x)+\max_{y\in D}d_{S^{\prime}}(1,y)$ and $C_{2}=\max_{s^{\prime}\in S^{\prime}}d_{S}(1,\phi^{\prime}(s^{\prime}))+\max_{s\in S}d_{S^{\prime}}(1,\phi(s))$. For any $g\in G$ with $d_{S}(1,g)=n$, it follows from Lemma 2.1 (1) that $\phi(g)\in\phi(S)^{n}D^{n-1}$. Then by triangle inequality, one gets that $d_{S^{\prime}}(1,\phi(g))\leq(C_{1}+C_{2})d_{S}(1,g)-C_{1}$. In general, for any $g_{1},g_{2}\in G$, Lemma 2.3 shows that there exists $a\in D^{-1}(D^{2}D^{-1})^{2}$ such that $\phi(g_{1})^{-1}\phi(g_{2})=\phi(g_{1}^{-1}g_{2})a$. Therefore,
$d_{S^{\prime}}(\phi(g_{1}),\phi(g_{2}))$ $=d_{S^{\prime}}(1,\phi(g_{1})^{-1}\phi(g_{2}))=d_{S^{\prime}}(1,\phi(g_{1}^{-1}g_{2})a)$
$\leq d_{S^{\prime}}(1,\phi(g_{1}^{-1}g_{2}))+7C_{1}\leq(C_{1}+C_{2})d_{S}(1,g_{1}^{-1}g_{2})+6C_{1}=(C_{1}+C_{2})d_{S}(g_{1},g_{2})+6C_{1}.$
As for the other direction, one has
$d_{S}(g_{1},g_{2})\leq d_{S}(\phi^{\prime}(\phi(g_{1})),\phi^{\prime}(\phi(g_{2})))+2C\leq(C_{1}+C_{2})d_{S^{\prime}}(\phi(g_{1}),\phi(g_{2}))+6C_{1}+2C.$
In summary, $\phi$ is a $(C_{1}+C_{2},6C_{1}+2C)$-quasi-isometry. ∎
The following lemma shows that those quasi-isomorphisms close to identity maps can be characterized by two special kinds of subsets in groups.
Lemma 2.11.
Let $\phi:G\to G$ be a map from a discrete group $G$ to itself. The following are equivalent:
  1. (1)
    $\phi$ is a quasi-isomorphism with a quasi-inverse given by the identity map.
  2. (2)
    The set $A=\{\phi(g)^{-1}g:g\in G\}$ is finite and $\phi(G)^{-1}$ almost commutes with $A$.
Proof.
[…]
$gh\sim_{A}\phi(gh)\sim_{D}\phi(g)\phi(h)\sim_{A^{-1}}\phi(g)h.$
By multiplying $h^{-1}\phi(g)^{-1}$ to the left, one gets that $\{h^{-1}\phi(g)^{-1}gh:g\in G,h\in\phi(G)\}\subseteq A^{-1}DA$. This implies that $\mathcal{C}(\phi(G)^{-1},A)$ is contained in $A^{-1}DAA^{-1}$ and thus is finite.
[…]
Note that for any $g\in G$, $g=\phi(g)a$ for some $a\in A$. Then for any $g=\phi(g)a,g^{\prime}=\phi(g^{\prime})a^{\prime}$, one has that
$\phi(gg^{\prime})\sim_{A^{-1}}gg^{\prime}=\phi(g)a\phi(g^{\prime})a^{\prime}=\phi(g)\phi(g^{\prime})\cdot[\phi(g^{\prime})^{-1},a]aa^{\prime}\sim_{\mathcal{C}(\phi(G)^{-1},A)A^{2}}\phi(g)\phi(g^{\prime}).$
This shows that $\phi$ is a quasi-homomorphism. ∎
3. Quasi-retractions
3.1. Elementary properties of quasi-retracts
Definition 3.1.
Let $(G,H)$ be a group pair. Classically a group homomorphism $r:G\to H$ is called a retraction if it is left-inverse to the inclusion $\iota:H\to G$. We thus call a quasi-homomorphism $r:G\to H$ a quasi-retraction if $\mathrm{dist}(r\circ\iota,Id_{H})<\infty$. If such a map exists, then $H$ is called a quasiretractquasi-retract of G.
At first, we give an equivalent definition of quasi-retraction.
Lemma 3.2.
Let $(G,H)$ be a group pair and $\iota:H\to G$ be the inclusion map. Then $H$ is a quasi-retract of $G$ if and only if there exists a quasi-homomorphism $r:G\to H$ such that $r\circ\iota=Id_{H}$.
Proof.
$\Leftarrow$” is obvious. We now prove $\Rightarrow$”. Let $\phi:G\to H$ be a quasi-retraction with the defect set $D=D(\phi)$. By definition, the set $A:=\{\phi(h)^{-1}h:h\in H\}$ is finite. Note that $h\sim_{A}\phi(h)$ for any $h\in H$.
Claim.
The set $B:=\{\phi(g)^{-1}a\phi(g):g\in G,a\in A\}$ is finite in $H$.
Proof of Claim.
For any $g\in G,h\in H$, one has $h\phi(g)\in H$. Thus,
$h\phi(g)\sim_{A}\phi(h\phi(g))\sim_{D}\phi(h)\phi(\phi(g))\sim_{A^{-1}}\phi(h)\phi(g).$
By multiplying $\phi(g)^{-1}\phi(h)^{-1}$ to the left, one gets that $B$ is contained in $A^{-1}DA$ and thus is finite. ∎
Define a map $r:G\to H$ as follows:
$r(g)=\left\{\begin{array}[]{ll}g,&\hbox{$g\in H$;}\\ \phi(g),&\hbox{$g\notin H$.}\end{array}\right$.
By definition, $r|_{H}=Id_{H}$. It remains to show that $r$ is a quasi-homomorphism, i.e. there exists a finite subset $D^{\prime}\subset H$ such that $r(gg^{\prime})\sim_{D^{\prime}}r(g)r(g^{\prime})$ for any $g,g^{\prime}\in G$.
Case I: $g,g^{\prime}\in H\Rightarrow gg^{\prime}\in H$. In this case, $r(gg^{\prime})=gg^{\prime}=r(g)r(g^{\prime})$.
Case II: $g\in H,g^{\prime}\notin H\Rightarrow gg^{\prime}\notin H$. In this case,
$r(g)r(g^{\prime})=g\phi(g^{\prime})=\phi(g)\phi(g^{\prime})\cdot\phi(g^{\prime})^{-1}\phi(g)^{-1}g\phi(g^{\prime})\sim_{B}\phi(g)\phi(g^{\prime})\sim_{D^{-1}}\phi(gg^{\prime})=r(gg^{\prime}).$
Case III: $g\notin H,g^{\prime}\in H\Rightarrow gg^{\prime}\notin H$. In this case,
$r(g)r(g^{\prime})=\phi(g)g^{\prime}\sim_{A}\phi(g)\phi(g^{\prime})\sim_{D^{-1}}\phi(gg^{\prime})=r(gg^{\prime}).$
Case IV: $g,g^{\prime}\notin H$. If $gg^{\prime}\in H$, then
$r(gg^{\prime})=gg^{\prime}=\phi(g)\phi(g^{\prime})\cdot\phi(g^{\prime})^{-1}\phi(g)^{-1}g\phi(g^{\prime})\cdot\phi(g^{\prime})^{-1}g^{\prime}\sim_{BA}\phi(g)\phi(g^{\prime})=r(g)r(g^{\prime}).$
If $gg^{\prime}\notin H$, then
$r(gg^{\prime})=\phi(gg^{\prime})\sim_{D}\phi(g)\phi(g^{\prime})=r(g)r(g^{\prime}).$
In summary, we conclude that $r$ is a quasi-homomorphism and the conclusion follows. ∎
With the help of Lemma 3.2, we will frequently use the equivalent definition of quasi-retracts in what follows.
Recall that a quasi-homomorphism with values in $\mathbb{R}$ equipped with the standard metric is called a quasimorphism. A quasimorphism $\phi:G\to\mathbb{R}$ is called homogeneous if $\phi(g^{n})=n\phi(g)$ for all $g\in G,n\in\mathbb{Z}$. According to [8, Lemma 2.21], any quasimorphism $\phi:G\to\mathbb{R}$ has a homogenization which is a homogeneous quasimorphism $\bar{\phi}:G\to\mathbb{R}$ defined by $\bar{\phi}(g)=\lim_{n\to\infty}\phi(g^{n})/n$.
Lemma 3.3.
A group $G$ admits an unbounded quasimorphism if and only if it contains an infinite cyclic quasi-retract.
Proof.
$\Leftarrow$” is clear. Now we prove $\Rightarrow$”. Let $\phi:G\to\mathbb{R}$ be an unbounded quasimorphism on $G$. Denote by $\bar{\phi}$ the homogenization of $\phi$. Since $\mathrm{dist}(\bar{\phi},\phi)<\infty$ (cf. [8, Lemma 2.21]), $\bar{\phi}\nequiv 0$. Up to rescaling $\bar{\phi}$, we can pick an element $h\in G$ such that $\bar{\phi}(h)=1$. Let $\lfloor\cdot\rfloor:\mathbb{R}\to\mathbb{Z}$ be the floor function which maps a real number $x$ to the maximal integer $\leq x$. Then the map $r:G\to\langle h\rangle$ defined by mapping each $g\in G$ to $r(g):=h^{\lfloor\bar{\phi}(g)\rfloor}$ gives a quasi-retraction. ∎
The next lemma gives some basic but very useful properties of quasi-retractions.
Lemma 3.4.
Let $r:G\to H$ be a quasi-retraction with $D=D(r)$. Then the set $A:=\{r(g)^{-1}g:g\in G\}$ satisfies $r(A)\subset D$. Moreover, suppose that $G$ is generated by a finite set $S$. Then
  1. (1)
    $H$ is generated by a finite set $r(S)\cup D$.
  2. (2)
    $r$ is a Lipschitz map with respect to $d_{S}$ on $G$ and any proper word metric on $H$.
  3. (3)
    $H$ is an undistorted subgroup of $G$.
  4. (4)
    if $G$ is of type $F_{n}$ (or $FP_{n}$), then so is $H$.
  5. (5)
    if $G:H]<\infty$, then $r$ is a quasi-isomorphism with a quasi-inverse given by the inclusion map $H\to G$.
Proof.
For any $g\in G$, $r(g)^{-1}\in H$. Then $r(r(g)^{-1}g)\sim_{D}r(r(g)^{-1})r(g)=r(g)^{-1}r(g)=1$.
(1) Let $S$ be a finite generating set of $G$. For any $h\in H\leq G$, we denote $h=s_{1}\cdots s_{n}$ where $s_{i}\in S,1\leq i\leq n$. Since $r|_{H}=Id_{H}$,
$h=r(h)=r(s_{1}\cdots s_{n})\sim_{D^{n-1}}r(s_{1})\cdots r(s_{n})\subset r(S)^{n}.$
Since $h$ is arbitrary, the above equality implies that $r(S)\cup D\subset H$ generates $H$.
(2) By Item (1), $S^{\prime}=r(S)\cup D$ is a finite generating set of $H$. Fix arbitrary two elements $g,g^{\prime}\in G$. Denote $n=d_{S}(g,g^{\prime})$. Namely, $g^{-1}g^{\prime}=s_{1}\cdots s_{n}$ where $s_{i}\in S,1\leq i\leq n$. Then
$r(g^{-1}g^{\prime})=r(s_{1}\cdots s_{n})\sim_{D^{n-1}}r(s_{1})\cdots r(s_{n})$
which implies that $d_{S^{\prime}}(1,r(g^{-1}g^{\prime}))\leq 2n-1$. Moreover, there exists $a\in D$ such that
$d_{S^{\prime}}(r(g),r(g^{\prime}))=d_{S^{\prime}}(r(g),r(g)r(g^{-1}g^{\prime})a)=d_{S^{\prime}}(1,r(g^{-1}g^{\prime})a)\leq 2n-1+1=2d_{S}(g,g^{\prime}).$
Since $g,g^{\prime}\in G$ are arbitrary, we obtain that $r$ is a Lipschitz map with respect to $d_{S}$ on $G$ and $d_{S^{\prime}}$ on $H$. The conclusion then follows since any two proper word metrics on $H$ are bi-Lipschitz to each other.
(3) Fix a finite generating set $S$ of $G$. By Item (1), $S^{\prime}=r(S)\cup D$ is a finite generating set of $H$. Let $S^{\prime\prime}=S\cup S^{\prime}$. Then $S^{\prime\prime}$ is also a finite generating set of $G$. Observe that $r(S^{\prime\prime})=S^{\prime}$. For the inclusion map $(H,d_{S^{\prime}})\to(G,d_{S^{\prime\prime}})$, the same proof of Item (2) shows that $d_{S^{\prime}}(h,h^{\prime})\leq 2d_{S^{\prime\prime}}(h,h^{\prime})$ for any $h,h^{\prime}\in H$. Since $d_{S^{\prime\prime}}(h,h^{\prime})\leq d_{S^{\prime}}(h,h^{\prime})$ is obvious, one gets that the inclusion map is a quasi-isometric embedding. Hence, $H$ is undistorted in $G$.
(4) This follows from the combination of Item (2) and [2, Theorem 8].
(5) Let $\iota:H\to G$ be the inclusion map. By definition of quasi-retracts, we have $r\circ\iota=Id_{H}$. It remains to show that $\mathrm{dist}(\iota\circ r,Id_{G})=\sup_{g\in G}d_{S}(r(g),g)<\infty$.
Fix a finite right coset transversal $T$ of $H$ in $G$. Then any $g\in G$ can be written uniquely as $g=g_{H}g_{T}$ where $g_{H}\in H$ and $g_{T}\in T$. Thus, $r(g)=r(g_{H}g_{T})=r(g_{H})r(g_{T})a=g_{H}r(g_{T})a$ for some $a\in D$. By setting $M:=\max_{x\in T^{-1}r(T)D}d_{S}(1,x)$, one obtains that
$d_{S}(r(g),g)=d_{S}(g_{H}r(g_{T})a,g_{H}g_{T})=d_{S}(1,g_{T}^{-1}r(g_{T})a)\leq M.$
Since $g$ is arbitrary, we complete the proof. ∎
Recall that a finitely generated group $G$ has property (QT) if $G$ acts isometrically on a finite product of quasi-trees equipped with $\ell^{1}$-metric such that the orbit map is a quasi-isometric embedding. This property generalizes the notion of finite asymptotic dimensions.
Corollary 3.5.
  1. (1)
    Quasi-retracts of hyperbolic groups are still hyperbolic groups.
  2. (2)
    Property (QT) is inherited by quasi-retracts.
Proof.
Both items follow directly from Lemma 3.4 (1) and (3). ∎
Since the coboundedness of group actions on hyperbolic spaces do not pass to undistorted subgroups, we raise the following question.
Question 2.
Is property (PH) or (PH’) inherited by quasi-retracts?
By the way, the following Corollary 5.10 shows that property (PH’) is inherited by normal quasi-retracts.
3.2. Quasi-retracts of hyperbolic groups
In this subsection, we are going to give a characterization of quasi-retracts of torsion-free hyperbolic groups and prove Theorem 1.1 (2). For an algebraic characterization of retracts of $\mathbb{F}_{n}$, we refer the readers to [31, Proposition 1].
Lemma 3.6.
Let $G$ be a torsion-free hyperbolic group and $H$ be a quasi-retract of $G$. Then $H$ is either a cyclic subgroup or a retract of $G$.
Proof.
Let $r:G\to H$ be a quasi-retraction and $D(r)$ be the defect set. By Corollary 3.5, $H$ is also a hyperbolic group. By Lemma 2.2, the defect subgroup $\Delta_{r}=\langle D(r)\rangle$ is a virtually abelian normal subgroup of $H$. Since a hyperbolic group does not contain $\mathbb{Z}^{2}$, the defect subgroup $\Delta_{r}$ must be cyclic. If $\Delta_{r}=\{1\}$, then $r$ is a retraction. Otherwise $\Delta_{r}$ is generated by a hyperbolic element. Then it follows from $\Delta_{r}\lhd H$ that $H$ is also cyclic. ∎
Remark 3.7.
If $G$ is a free group in Lemma 3.6, then every cyclic subgroup is also a quasi-retract of $G$. In fact, suppose that $H=\langle w\rangle$ for some $w\in G$. Up to conjugation, we can assume that $w$ is a cyclically reduced word in $G$. Let $h_{w}$ be the corresponding Brooks counting quasimorphism, which is a quasimorphism on $G$ satisfying $h_{w}(w^{n})=n$ for $n\in\mathbb{Z}$; see [8, Subsection 2.3.3] for a reference. Then the map $r:G\to H$ defined by $r(g)=w^{h_{w}(g)}$ is a quasi-retraction.
Recall that a quasi-retract is called nontrivial if it is neither a finite subgroup nor the whole group. A group element $g\in G$ is called primitive if $g$ can not be represented as a nonzero power of any $h\in G\setminus\{g,g^{-1}\}$. A group $G$ is called Hopfian if any surjective homomorphism from $G$ to $G$ is an isomorphism.
Lemma 3.8.
Let $G=\mathbb{F}_{2}$. Any nontrivial retract of $G$ is a cyclic subgroup generated by a primitive element.
Proof.
Let $H\leq G$ be a nontrivial retract of $G$. That is, there exists a homomorphism $r:G\to H$ such that $r|_{H}=Id$. Since $r$ is a surjective homomorphism, $\mathrm{rk}(H)\leq\mathrm{rk}(G)=2$. Since $H$ is nontrivial, $\mathrm{rk}(H)\geq 1$.
If $\mathrm{rk}(H)=2$, then $H\cong\mathbb{F}_{2}$ since any subgroup of a free group is also free. Another well-known fact is that any free group of finite rank is Hopfian [27, Theorem 3.3]. This implies that $r$ is an isomorphism, which contradicts with that $H\neq G$. Therefore $\mathrm{rk}(H)=1$.
Let $H=\langle g\rangle$. Suppose that $g$ is not primitive, i.e. $g=h^{n}$ for some $h\neq 1\in G,n\geq 2$. Denote $r(h)=h^{k}$ for some $k\in\mathbb{Z}$. Then $h^{n}=g=r(g)=r(h^{n})=r(h)^{n}=h^{kn}$ which implies that $k=1$. However, $r(h)=h\notin H$ which is impossible. Therefore, $g$ is a primitive element. ∎
A subgroup $H\leq G$ is called malnormal if $H\cap gHg^{-1}=\{1\}$ for any $g\in G\setminus H$.
Proposition 3.9.
A subgroup of $\mathbb{F}_{2}$ is a nontrivial quasi-retract if and only if it is a cyclic group. Moreover, there exists a hyperbolically embedded subgroup of $\mathbb{F}_{2}$ which is not a quasi-retract.
Proof.
If $H$ is a nontrivial quasi-retract of $\mathbb{F}_{2}$, then Lemma 3.6, together with Lemma 3.8 show that $H$ is a cyclic group. The opposite direction is given by Remark 3.7. Now we prove the “Moreover” part. In [4, Example 1], Baumslag-Miasnikov-Remeslennikov constructed an example of a finitely generated malnormal subgroup $H$ of $\mathbb{F}_{2}$ which has rank greater than 2. Hence, it can not be a quasi-retract of $\mathbb{F}_{2}$. Since any finitely generated subgroup of a free group is quasi-convex, it follows from [6] that $\mathbb{F}_{2}$ is hyperbolic relative to $H$. As a result of [9, Proposition 2.4], $H$ is a hyperbolically embedded subgroup of $\mathbb{F}_{2}$. ∎
3.3. First examples
The goal of this subsection is to give some general examples and non-examples of quasi-retracts. The first part aims to explore when a finite-index subgroup becomes a quasi-retract. Note that a quasimorphism on a finite-index subgroup can not be extended to the whole group in general.
Proposition 3.10.
Let $(G,H)$ be a group pair. Suppose $H$ is of finite index in $G$. Then $H$ is a quasi-retract of $G$ if and only if there exists a right coset transversal $T$ containing 1 of $H$ in $G$ such that $H$ almost commute with $T$.
Proof.
$\Rightarrow$”: let $r:G\to H$ be a quasi-retraction with $D=D(r)$. Since $G:H]<\infty$, Lemma 3.4 (5) shows that $r$ is a quasi-isomorphism with a quasi-inverse given by the inclusion map $H\to G$. As a result of Lemma 2.11, the set $A:=\{r(g)^{-1}g:g\in G\}$ is finite and $H$ almost commutes with $A$. Note that any $g\in G$ can be written as $g=r(g)\cdot r(g)^{-1}g$ and $1\in A$. Hence, $A$ contains a right coset transversal $T$ containing $1$ of $H$ in $G$. And the finiteness of $\mathcal{C}(H,T)$ follows from the finiteness of $\mathcal{C}(H,A)$.
$\Leftarrow$”: suppose that there exists a right coset transversal $T$ containing 1 of $H$ in $G$ such that $H$ almost commute with $T$. For any $\in G, gg$ can be uniquely written as $g=g_{H}g_{T}$ where $g_{H}\in H$ and $g_{T}\in T$. Define a map $r:G\to H,g=g_{H}g_{T}\mapsto g_{H}$. Since $1\in T$, $r|_{H}=Id_{H}$. Since $T$ is finite, the set $A:=\{r(g)^{-1}g:g\in G\}\subseteq T$ is also finite. By Lemma 2.11, $r$ is a quasi-homomorphism (even a quasi-isomorphism) and the conclusion then follows. ∎
We also show that a quasi-retraction to a finite-index subgroup of a subgroup is a composition of two quasi-retractions.
Lemma 3.11.
Let $(G,H)$ be a group pair. Let $K$ be a finite-index subgroup of $H$. Then $K$ is a quasi-retract of $G$ if and only if $H$ is a quasi-retract of $G$ and $K$ is a quasi-retract of $H$.
Proof.
$\Leftarrow$” is clear since a composition of two quasi-retractions is still a quasi-retraction. Now, we prove $\Rightarrow$”. Let $r:G\to K$ be a quasi-retraction with the defect set $D=D(r)$. Clearly $r|_{H}:H\to K$ is also a quasi-retraction. Since $H:K]<\infty$, Lemma 3.4 (5) shows that $r|_{H}$ is a quasi-isomorphism with a quasi-inverse given by the inclusion map $K\to H$. By Lemma 2.11, the set $A:=\{r(h)^{-1}h:h\in H\}$ is finite. Now, we take $r:G\to K$ as a quasi-homomorphism $r:G\to H$. The finiteness of $A$ implies that $\mathrm{dist}(r|_{H},Id_{H})<\infty$. By definition, $H$ is a quasi-retract of $G$. ∎
Before giving more general examples of quasi-retracts, we first recall some definitions.
Definition 3.12.
[25] A group $G$ is said to be the direct product of its subgroups $H$ and $K$ with an amalgamated subgroup $Z$, denoted by $H\times_{Z}K$ if the following conditions hold: (i) $G$ is generated by $H\cup K$, (ii) $H\cap K=Z$, (iii) $H$ commutes with $K$, i.e. $\mathcal{C}(H,K)=\{1\}$.
By [12, Lemma 2.9], a central extension of groups $1\to H\to G\to Q\to 1$ has a bounded Euler class if and only if the central extension is right quasi-split, i.e. there exists a section $s:Q\to G$ such that $s$ is a quasi-homomorphism.
Definition 3.13.
[30, Definition 2.1] A group $G$ has property (PH’) if there exist finitely many hyperbolic spaces $X_{1},\ldots,X_{n}$ on which $G$ acts coboundedly and the induced diagonal action of $G$ on the product $\prod_{i=1}^{n}X_{i}$ with $\ell^{1}$-metric is proper. In this case, we also say $G$ has property (PH’) from actions $G\curvearrowright X_{1},\cdots,G\curvearrowright X_{n}$.
A group $G$ has property (PH) if $G$ virtually has property (PH’).
By definition, any finitely generated abelian group has property (PH).
Lemma 3.14.
Let $(G,H)$ be a group pair. Then $H$ is a quasi-retract of $G$ whenever the pair satisfies one of the following:
  1. (1)
    $H$ is a retract of $G$. Equivalently, $G$ has a semi-direct product structure of the form $G=K\rtimes H$.
  2. (2)
    $G$ is an arbitrary group and $H$ is a finite subgroup.
  3. (3)
    $G$ is an amalgamated direct product of the form $H\times_{Z}K$ and the central extension $1\to Z\to K\to K/Z\to 1$ has a bounded Euler class.
  4. (4)
    $G$ has property (PH) and $H\leq Z(G)$.
  5. (5)
    $G$ is a finitely generated abelian group and $H$ is an arbitrary subgroup.
  6. (6)
    $H\leq Z(G)$ and the central extension $1\to H\to G\to G/H\to 1$ has a bounded Euler class.
Proof.
(1) is clear since a homomorphism is obviously a quasi-homomorphism. (2) is clear by mapping every element in $H$ to itself and every element outside $H$ to $1$. (6) is given by [30, Lemma 5.6]. By [30, Theorem 1.7], (4) is transformed into (6). (5) follows directly from (4). Finally, it remains to show (3). Since the central extension $1\to Z\to K\to K/Z\to 1$ has a bounded Euler class, there exists a section $s:K/Z\to K$ such that $s$ is a quasi-homomorphism. By Lemma 2.6, we can assume that this section is normalized, i.e. $s(1)=1$. Fix a right coset transversal $T$ containing $1$ of $Z$ in $H$. By [18, Corollary 6], any element $g\in H\times_{Z}K$ can be written uniquely as $g=hkz$ where $h\in T,k\in s(K/Z),z\in Z$. Define a map $r:G\to H$ by sending $g=hkz\mapsto hz$. Clearly $r|_{H}=Id_{H}$. Since $H$ commutes with $K$, it is direct to compute that $D(r)\subseteq D(s)^{-1}$. Hence, $r$ is a quasi-homomorphism and then quasi-retraction.∎
We also collect some concrete non-examples of quasi-retractions.
Example 3.15.
  1. (1)
    Let $G=\langle t,s\mid s^{2}=(st)^{2}=1\rangle$ be an infinite dihedral group and $H=\langle t\rangle$. Since $[G:H]=2$ and $H$ does not almost commute with any right coset transversal $T$ of $H$ in $G$, Proposition 3.10 implies that $H$ is not a quasi-retract of $G$.
  2. (2)
    Let $G=\langle a,b,c\mid[a,c]=[b,c]=1,[a,b]=c\rangle$ be the Heisenberg group and $H=\langle c\rangle=Z(G)$. Since $H$ is distorted in $G$, Lemma 3.4 (3) implies that $H$ is not a quasi-retract of $G$. However, $G$ quasi-retracts to $\langle a\rangle$ since $G$ has a homomorphism to $\langle a\rangle$ by sending $a,b,c$ to $a,1,1$ respectively.
  3. (3)
    Let $G=BS(1,n)=\langle a,t\mid tat^{-1}=a^{n}\rangle,n\neq 0,1$ be a Baumslag-Solitar group and $H=\langle a\rangle$. Since $H$ is distorted in $G$, Lemma 3.4 (3) implies that $H$ is not a quasi-retract of $G$. However, $G$ quasi-retracts to $\langle t\rangle$ since $G$ has a homomorphism to $\langle t\rangle$ by sending $t,a$ to $t,1$ respectively.
4. Quasi-split short exact sequences
Recall from the Introduction that, a short exact sequence $1\to H\to G\to Q\to 1$ is left split if $H$ is a retract of $G$ and right split if it admits a section $s:Q\to G$ that is a homomorphism. Comparing with the classical notion, we say the short exact sequence is left quasi-split if $H$ is a quasi-retract of $G$ and right quasi-split if it admits a section $s:Q\to G$ that is a quasi-homomorphism. One simple observation is that whenever $H$ or $Q$ is finite, the above short exact sequence is always right quasi-split since the defect set of any section is a finite subset of $H$.
Recall in the classical group theory, we have the following results.
Theorem 4.1 (Classical version).
Let $1\to H\xrightarrow{\iota}G\xrightarrow{\pi}Q\to 1$ be a short exact sequence of groups. The following are equivalent:
  1. (1)
    The short exact sequence $1\to H\to G\to Q\to 1$ is left split. Or equivalently, $H$ is a retract of $G$.
  2. (2)
    There exists a section $s:Q\to G$ such that $s$ is a homomorphism and $H$ commutes with $s(Q)$.
  3. (3)
    There is an isomorphism $\phi:G\to H\times Q$ such that the diagram commutes:
[…]
    where the bottom row is the short exact sequence for a direct product.
Before giving the analogue of Theorem 4.1, we need the following two lemmas. Recall that for any two subsets $H,T$ of a group $G$, we denote $\mathcal{C}(H,T):=\{[h,t]=hth^{-1}t^{-1}:h\in H,t\in T\}$.
Lemma 4.2.
Let $G$ be a group and $H\lhd G$ be a normal subgroup. For any subset $T\subset G$, $\mathcal{C}(H,T)=\mathcal{C}(T^{-1},H)=\mathcal{C}(H,T^{-1})^{-1}$.
Proof.
Let $T\subset G$ be a subset. Since $t^{-1},h]=t^{-1}hth^{-1}=(ht^{-1}h^{-1}t)^{-1}=[h,t^{-1}]^{-1}$ for all $h\in H,t\in T$, it is clear that $\mathcal{C}(T^{-1},H)=\mathcal{C}(H,T^{-1})^{-1}$. Thus we only need to verify that $\mathcal{C}(H,T)=\mathcal{C}(T^{-1},H)$. Since $H\lhd G$, $tht^{-1},t^{-1}ht\in H$ for any $h\in H,t\in T$. Denote $h^{t}:=tht^{-1}$ and $h^{t^{-1}}=t^{-1}ht$. Then $(h^{-1})^{t}=th^{-1}t^{-1}=(h^{t})^{-1}$ and $(h^{-1})^{t^{-1}}=t^{-1}h^{-1}t=(h^{t^{-1}})^{-1}$. Therefore,
$h,t]=hth^{-1}t^{-1}=t^{-1}(tht^{-1})t(th^{-1}t^{-1})=t^{-1}h^{t}t(h^{-1})^{t}=[t^{-1},h^{t}]\in\mathcal{C}(T^{-1},H)$
and similarly
$t^{-1},h]=t^{-1}hth^{-1}=(t^{-1}ht)t(t^{-1}h^{-1}t)t^{-1}=h^{t^{-1}}t(h^{-1})^{t^{-1}}t^{-1}=[h^{t^{-1}},t]\in\mathcal{C}(H,T).$
Since $h\in H,t\in T$ are arbitrary, the above formulas show that $\mathcal{C}(H,T)\subset\mathcal{C}(T^{-1},H)\subset\mathcal{C}(H,T)$. Thus we get $\mathcal{C}(H,T)=\mathcal{C}(T^{-1},H)$ and the conclusion follows. ∎
The following lemma gives a useful necessary condition when a subgroup is a quasi-retract.
Lemma 4.3.
Let $(G,H)$ be a group pair. If $r:G\to H$ is a quasi-retraction with $D=D(r)$, then for any normal subgroup $H_{0}\lhd G$ contained in $H$, there exists a right coset transversal $T$ containing $1$ of $H$ in $G$ such that $r(T)\subset D$ and $\mathcal{C}(H_{0},T)$ is finite.
Proof.
Since $r(1)=1=r(gg^{-1})\sim_{D}r(g)r(g^{-1})$, one has that $r(g^{-1})\sim_{D^{-1}}r(g)^{-1}$ for any $g\in G$. Let $T$ be an arbitrary right coset transversal containing $1$ of $H$ in $G$. For each $t\in T$, we denote $t^{\prime}:=r(t)^{-1}t$ and $T^{\prime}=\{t^{\prime}:t\in T\}$. Clearly $T^{\prime}$ gives another right coset transversal containing $1$ of $H$ in $G$. And Lemma 3.4 shows that $r(T^{\prime})\subset D$.
Since $H_{0}\lhd G$ and $r|_{H_{0}}=Id_{H_{0}}$, one has that for any $h\in H_{0},t^{\prime}\in T^{\prime}$,
$t^{\prime}ht^{\prime-1}=r(t^{\prime}ht^{\prime-1})\sim_{D^{2}}r(t^{\prime})r(h)r(t^{\prime-1})=r(t^{\prime})hr(t^{\prime-1})\sim_{D^{-1}}r(t^{\prime})hr(t^{\prime})^{-1}\sim_{D^{-1}}r(t^{\prime})h.$
By multiplying $h^{-1}r(t^{\prime})^{-1}$ to the left, the above formula transforms into
$h^{-1}r(t^{\prime})^{-1}t^{\prime}ht^{\prime-1}\sim_{D^{-2}D^{2}}1.$
Denote $\tilde{t}:=r(t^{\prime})^{-1}t^{\prime}$. Therefore, we get
$h^{-1},\tilde{t}]=h^{-1}\tilde{t}h\tilde{t}^{-1}=h^{-1}r(t^{\prime})^{-1}t^{\prime}ht^{\prime-1}r(t^{\prime})\sim_{D}h^{-1}r(t^{\prime})^{-1}t^{\prime}ht^{\prime-1}\sim_{D^{-2}D^{2}}1.$
Since $h\in H_{0}$ is arbitrary and $\tilde{T}=\{\tilde{t}:t\in T\}$ is still a right coset transversal containing $1$ of $H$ in $G$, the conclusion follows. ∎
Remark 4.4.
We remark that the inverse of Lemma 4.3 is not true. For example, let $G$ be a Heisenberg group and $H$ be its center. Then $\mathcal{C}(H_{0},T)=\{1\}$ since $H_{0}\leq H=Z(G)$. However, Example 3.15 (2) shows that $H$ is not a quasi-retract of $G$.
Recall that a quasi-homomorphism $\phi:G\to G^{\prime}$ is quasi-isomorphism if there exists a quasi-homomorphism $\phi^{\prime}:G^{\prime}\to G$, called the quasi-inverse of $\phi$, such that $\mathrm{dist}(\phi^{\prime}\circ\phi,Id_{G})<\infty$ and $\mathrm{dist}(\phi\circ\phi^{\prime},Id_{G^{\prime}})<\infty$. Moreover, if $\phi^{\prime}=\phi^{-1}$, then $\phi$ is a strict quasi-isomorphism.
Theorem 4.5.
Let $1\to H\xrightarrow{\iota}G\xrightarrow{\pi}Q\to 1$ be a short exact sequence of groups. The following are equivalent:
  1. (1)
    The short exact sequence $1\to H\to G\to Q\to 1$ is left quasi-split. Or equivalently, $H$ is a quasi-retract of $G$.
  2. (2)
    There exists a normalized section $s:Q\to G$ such that $s$ is a quasi-homomorphism and $H$ almost commutes with $s(Q)$: $\mathcal{C}(H,s(Q))$ is finite.
  3. (3)
    There is a strict quasi-isomorphism $\phi:G\to H\times Q$ such that the diagram commutes:
[…]
    where the bottom row is the short exact sequence for a direct product.
Proof.
[…]
For any $g=g_{H}g_{T},g^{\prime}=g^{\prime}_{H}g^{\prime}_{T}\in G$, one has that $gg^{\prime}=g_{H}g_{T}g^{\prime}_{H}g^{\prime}_{T}=(gg^{\prime})_{H}(gg^{\prime})_{T}$. Then Lemma 2.7 shows that
$r(gg^{\prime})=(gg^{\prime})_{H}=g_{H}(g_{T}g^{\prime}_{H}g_{T}^{-1})(g_{T}g^{\prime}_{T})_{H}=g_{H}g^{\prime}_{H}\cdot[{g^{\prime}_{H}}^{-1},g_{T}](g_{T}g^{\prime}_{T})_{H}\sim_{\mathcal{C}(H,T)\tilde{D}(s)}g_{H}g^{\prime}_{H}=r(g)r(g^{\prime}).$
The above equality shows that the defect set $D(r)$ of $r$ is contained in $\mathcal{C}(H,T)\tilde{D}(s)$ and thus is finite.
[…]
It remains to verify that $s$ is a quasi-homomorphism. Note that for any $g\in G$, $r(g^{-1})\sim_{D^{-1}}r(g)^{-1}$ and for any $x,y\in Q$, $r(s(x)),r(s(y)),r(s(xy))\in r(T)\subset D$. Since $s$ is a section, $s(y)^{-1}s(x)^{-1}s(xy)\in H$. Therefore,
$s(y)^{-1}s(x)^{-1}s(xy)$ $=r(s(y)^{-1}s(x)^{-1}s(xy))\sim_{D^{2}}r(s(y)^{-1})r(s(x)^{-1})r(s(xy))$
$\sim_{D}r(s(y)^{-1})r(s(x)^{-1})\sim_{D^{-1}}r(s(y)^{-1})r(s(x))^{-1}\sim_{D^{-1}}r(s(y)^{-1})\sim_{D^{-2}}1.$
Since $x,y\in Q$ are arbitrary, the above formula shows that the defect of $s$ is finite, which means that $s$ is a quasi-homomorphism.
[…]
$\phi(g)=(r(g),\pi(g)),\quad\phi^{\prime}(h,x)=h\cdot s(x).$
Since $r$ is a quasi-retraction and $\pi$ is a homomorphism, it is clear that $\phi$ is a quasi-homomorphism. As for $\phi^{\prime}$, we compute directly that
$\phi^{\prime}(h_{2},x_{2})^{-1}\phi^{\prime}(h_{1},x_{1})^{-1}\phi^{\prime}(h_{1}h_{2},x_{1}x_{2})=s(x_{2})^{-1}h_{2}^{-1}s(x_{1})^{-1}h_{2}s(x_{1}x_{2})$
$=s(x_{2})^{-1}[h_{2}^{-1},s(x_{1})^{-1}]s(x_{2})s(x_{2})^{-1}s(x_{1})^{-1}s(x_{1}x_{2})\sim_{D(s)}s(x_{2})^{-1}[h_{2}^{-1},s(x_{1})^{-1}]s(x_{2})$
$=[s(x_{2})^{-1},[h_{2}^{-1},s(x_{1})^{-1}]][h_{2}^{-1},s(x_{1})^{-1}]\sim_{\mathcal{C}(H,T^{-1})}[s(x_{2})^{-1},[h_{2}^{-1},s(x_{1})^{-1}]]\sim_{\mathcal{C}(T^{-1},H)}1$
for any $(h_{1},x_{1}),(h_{2},x_{2})\in H\times Q$. By Lemma 4.2, $\mathcal{C}(H,T^{-1})$ and $\mathcal{C}(T^{-1},H)$ are both finite since $\mathcal{C}(H,T)$ is finite. Therefore, $\phi^{\prime}$ is also a quasi-homomorphism.
It remains to show that $\phi^{\prime}=\phi^{-1}$. For any $g\in G$, $\phi^{\prime}(\phi(g))=\phi^{\prime}(r(g),\pi(g))=r(g)s(\pi(g))=g_{H}g_{T}=g$. For any $(h,x)\in H\times Q$, $\phi(\phi^{\prime}(h,x))=(r(hs(x)),\pi(hs(x)))=(h,x)$. Thus $\phi^{\prime}=\phi^{-1}$, which completes the proof.
[…]
Remark 4.6.
In the classical Theorem 4.1, the direct product structure of $G$ guarantees that there exists a section $s:Q\to G$ which is a homomorphism. However, different to the classical version, the finiteness of $\mathcal{C}(H,s(Q))$ does not imply that $s$ is a quasi-homomorphism in Theorem 4.5. For example, let $G$ be the discrete Heisenberg group and $H=Z(G)$. Then $\mathcal{C}(H,s(G/H))=\{1\}$ for any normalized section $s:G/H\to G$. However, Example 3.15 (2) shows that the central extension $1\to H\to G\to G/H\to 1$ is not left quasi-split. Thus as a result of Theorem 4.5, any normalized section $s:G/H\to G$ can not be a quasi-homomorphism. By the way, Example 3.15 (1) shows that one can not drop the requirement that $\mathcal{C}(H,s(Q))$ is finite in Theorem 4.5, either.
The following result is a complement to Lemma 3.6 which initially characterizes quasi-retracts of torsion-free hyperbolic groups.
Proposition 4.7.
  1. (1)
    A normal quasi-retract of a hyperbolic group is either finite or of finite index.
  2. (2)
    A normal quasi-retract of a non-elementary torsion-free hyperbolic group is trivial.
Proof.
(1) Let $G$ be a finitely generated hyperbolic group and $H\lhd G$ be a quasi-retract. By Lemma 3.4 (1), $H$ is finitely generated. By Theorem 4.5, $G$ is strictly quasi-isomorphic to $H\times G/H$. By Lemma 2.10, $G$ is quasi-isometric to $H\times G/H$. Since word-hyperbolicity is a quasi-isometry invariant for finitely generated groups. we know that $H\times G/H$ is also hyperbolic. A well-known fact is that a hyperbolic group can contain neither an infinite torsion subgroup nor $\mathbb{Z}^{2}$. This forces either $H$ is finite or $G/H$ is finite. Then the conclusion follows.
(2) Let $G$ be a finitely generated non-elementary torsion-free hyperbolic group and $H\lhd G$ be a quasi-retract. By Item (1), it suffices to show that $G/H=\{1\}$ if $G:H]<\infty$. By Theorem 4.5, there exists a normalized section $s:G/H\to G$ such that $H$ almost commutes with $s(G/H)$.
For each hyperbolic element $f\in G$, we denote $E(f):=\{g\in G:\exists\ n\geq 1,\text{ s.t. }gf^{n}g^{-1}=f^{n}\}$. It is well-known that in a non-elementary torsion-free hyperbolic group $G$, $E(f)$ is infinite cyclic for each hyperbolic element $f\in G$ and
$\bigcap_{f:\text{ hyperbolic in }G}E(f^{k})=\{1\}$
for each $k\geq 1$.
Since $G:H]<\infty$, there exists $k\geq 1$ such that $g^{k}\in H$ for all $g\in G$. Fix an arbitrary hyperbolic element $f\in G$. Then $f^{k}\in H$. It follows from the finiteness of $\mathcal{C}(H,s(G/H))$ that $\mathcal{C}(\langle f^{k}\rangle,s(G/H))$ is finite. This implies that $s(G/H)\subset E(f^{k})$. Since $f\in G$ is arbitrary, one gets that $s(G/H)\subseteq\bigcap_{f:\text{ hyperbolic in }G}E(f^{k})=\{1\}$. Then the conclusion follows. ∎
Let $(G,H)$ be a group pair. Recall from [9, Definition 2.1] that $H$ is hyperbolically embedded in $G$ if there exists a subset $X\subseteq G$ such that the following conditions hold:
  • (i)
    $G$ is generated by $X\cup H$.
  • (ii)
    The Cayley graph $\mathcal{G}(G,X\cup H)$ is hyperbolic.
  • (iii)
[…]
Proposition 4.8.
A finite-by-$\mathbb{Z}^{m}$ hyperbolically embedded subgroup is a quasi-retract.
Proof.
Let $G$ be a group and $H\leq G$ be a finite-by-$\mathbb{Z}^{m}$ hyperbolically embedded subgroup. By assumption, $H$ fits into the following short exact sequence of groups:
(3) $1\to F\to H\xrightarrow{\pi}\mathbb{Z}^{m}\to 1$
where $F\leq H$ is a finite group. Since a finite subgroup is always a quasi-retract, the above short exact sequence (3) is left quasi-split. By Theorem 4.5, there exists a normalized section $s:\mathbb{Z}^{m}\to H$ such that $s$ is a quasi-homomorphism and $\mathcal{C}(F,s(\mathbb{Z}^{m}))\subset F$ is finite. If $m=0$, then $H$ is a finite subgroup which is obviously a quasi-retract. Now we assume $m\geq 1$. Let $\bar{c}_{1},\cdots,\bar{c}_{m}$ be a standard generating set of $\mathbb{Z}^{m}$ and denote $c_{i}=s(\bar{c}_{i})$ for $1\leq i\leq m$. Then any $g\in H$ can be uniquely represented as $g=fc_{1}^{n_{1}}\cdots c_{m}^{n_{m}}$ for some $f\in F,n_{1},\cdots,n_{m}\in\mathbb{Z}$.
Fix $i\in\{1,\cdots,m\}$. The projection map $\phi_{i}:H\to\mathbb{Z},\quad g=fc_{1}^{n_{1}}\cdots c_{m}^{n_{m}}\mapsto n_{i}$ factors through $H\xrightarrow{\pi}\mathbb{Z}^{m}\to\mathbb{Z}$ and thus is a group homomorphism from $H$ to $\mathbb{Z}$. Since $H$ is hyperbolically embedded in $G$, the result of Hull-Osin (i.e. [17, Theorem 1.4]) shows that the homomorphism $\phi_{i}$ can be extended to a quasimorphism $\tilde{\phi}_{i}:G\to\mathbb{R}$ satisfying $\tilde{\phi}_{i}|_{H}=\phi_{i}$. Let $\lfloor\cdot\rfloor:\mathbb{R}\to\mathbb{Z}$ be the floor function which maps a real number $x$ to the maximal integer $\leq x$. Clearly $\lfloor\cdot\rfloor$ is a quasimorphism and so is the composite map $\rho_{i}=\lfloor\cdot\rfloor\circ\tilde{\phi}_{i}$. Note that $\rho_{i}|_{H}=\phi_{i}$. This implies that $h\sim_{F}c_{1}^{\rho_{1}(h)}\cdots c_{m}^{\rho_{m}(h)}$ for any $h\in H$.
Define a map $r:G\to H$ as follows:
$r(g)=\left\{\begin{array}[]{ll}g,&\hbox{$g\in H$;}\\ c_{1}^{\rho_{1}(g)}\cdots c_{m}^{\rho_{m}(g)},&\hbox{$g\notin H$.}\end{array}\right$.
By definition, $r|_{H}=Id_{H}$. It suffices to show that there exists a finite subset $D\subset H$ such that $r(gg^{\prime})\sim_{D}r(g)r(g^{\prime})$ for any $g,g^{\prime}\in G$. Let $D(\rho_{i})$ be the defect set of $\rho_{i}$ for $1\leq i\leq m$. Hence, the set $A:=\{c_{1}^{d_{1}}\cdots c_{m}^{d_{m}}:d_{1}\in D(\rho_{1}),\cdots,d_{m}\in D(\rho_{m})\}$ is finite in $H$. Also observe that $h_{1}\sim_{F}h_{2}$ whenever $h_{1},h_{2}\in H$ satisfies $\pi(h_{1})=\pi(h_{2})$.
Case I: $g,g^{\prime}\in H\Rightarrow gg^{\prime}\in H$. In this case, $r(gg^{\prime})=gg^{\prime}=r(g)r(g^{\prime})$.
Case II: $g\in H,g^{\prime}\notin H\Rightarrow gg^{\prime}\notin H$. In this case,
$r(gg^{\prime})=c_{1}^{\rho_{1}(gg^{\prime})}\cdots c_{m}^{\rho_{m}(gg^{\prime})}\sim_{A}c_{1}^{\rho_{1}(g)+\rho_{1}(g^{\prime})}\cdots c_{m}^{\rho_{m}(g)+\rho_{m}(g^{\prime})}\sim_{F}c_{1}^{\rho_{1}(g)}\cdots c_{m}^{\rho_{m}(g)}r(g^{\prime})\sim_{F}r(g)r(g^{\prime}).$
Case III: $g\notin H,g^{\prime}\in H\Rightarrow gg^{\prime}\notin H$. In this case,
$r(gg^{\prime})=c_{1}^{\rho_{1}(gg^{\prime})}\cdots c_{m}^{\rho_{m}(gg^{\prime})}\sim_{A}c_{1}^{\rho_{1}(g)+\rho_{1}(g^{\prime})}\cdots c_{m}^{\rho_{m}(g)+\rho_{m}(g^{\prime})}\sim_{F}r(g)c_{1}^{\rho_{1}(g^{\prime})}\cdots c_{m}^{\rho_{m}(g^{\prime})}\sim_{F}r(g)r(g^{\prime}).$
Case IV: $g,g^{\prime}\notin H$. If $gg^{\prime}\in H$, then
$r(gg^{\prime})=gg^{\prime}\sim_{F}c_{1}^{\rho_{1}(gg^{\prime})}\cdots c_{m}^{\rho_{m}(gg^{\prime})}\sim_{A}c_{1}^{\rho_{1}(g)+\rho_{1}(g^{\prime})}\cdots c_{m}^{\rho_{m}(g)+\rho_{m}(g^{\prime})}\sim_{F}r(g)r(g^{\prime}).$
If $gg^{\prime}\notin H$, then
$r(gg^{\prime})=c_{1}^{\rho_{1}(gg^{\prime})}\cdots c_{m}^{\rho_{m}(gg^{\prime})}\sim_{A}c_{1}^{\rho_{1}(g)+\rho_{1}(g^{\prime})}\cdots c_{m}^{\rho_{m}(g)+\rho_{m}(g^{\prime})}\sim_{F}r(g)r(g^{\prime}).$
In summary, we conclude that $r$ is a quasi-homomorphism and the conclusion follows. ∎
By combining Proposition 3.9 and Proposition 4.8, we complete the proof of Theorem 1.1.
5. Induced group actions on hyperbolic spaces
5.1. Preliminaries of group actions on hyperbolic spaces
For an isometric action of a group $G$ on a hyperbolic space $X$, we denote by $\partial X$ the Gromov boundary of $X$ and $\Lambda(G)$ the limit set of $G$ in $\partial X$.
By Gromov [14], the isometries of a hyperbolic space $X$ can be subdivided into three classes. A nontrivial element $g\in\mathrm{Isom}(X)$ is called elliptic if some $\langle g\rangle$-orbit is bounded. Otherwise, it is called hyperbolic (resp. parabolic) if it has exactly two fixed points (resp. one fixed point) in the Gromov boundary $\partial X$ of $X$.
Theorem 5.1.
[1, Theorem 4.2] Let $G$ be a group acting isometrically on a hyperbolic space $X$. Then exactly one of the following conditions holds.
  1. (1)
    $|\Lambda(G)|=0$. Equivalently, $G$ has bounded orbits. In this case the action of $G$ is called elliptic.
  2. (2)
    $|\Lambda(G)|=1$. Equivalently, $G$ has unbounded orbits and contains no hyperbolic elements. In this case the action of $G$ is called parabolic or horocyclic. A parabolic action cannot be cobounded and the set of points of $\partial X$ fixed by $G$ coincides with $\Lambda(G)$.
  3. (3)
    $|\Lambda(G)|=2$. Equivalently, $G$ contains a hyperbolic element and any two hyperbolic elements have the same fixed points in $\partial X$. In this case the action of $G$ is called lineal.
  4. (4)
    $|\Lambda(G)|=\infty$. Then $G$ always contains hyperbolic elements. In turn, this case breaks into two subcases.
    1. (a)
      $G$ fixes a point of $\partial X$. Equivalently, any two hyperbolic elements of $G$ have a common fixed point in the boundary. In this case the action of $G$ is called quasi-parabolic or focal. Orbits of quasi-parabolic actions are always quasi-convex.
    2. (b)
      $G$ does not fix any point of $\partial X$. Equivalently, $G$ contains infinitely many independent hyperbolic elements. In this case the action of $G$ is said to be of general type.
A quasi-line is a metric space quasi-isometric to $\mathbb{R}$ equipped with the standard metric. We list some well-known facts for future use.
Fact 1.
  1. (1)
    Any finitely generated group admitting a focal or general type action on a hyperbolic space contains a non-abelian free sub-semigroup and thus has exponential growth.
  2. (2)
    Group actions on quasi-lines are either elliptic or lineal.
  3. (3)
    If a group admits a cobounded lineal action on a hyperbolic space $X$, then $X$ is a quasi-line.
  4. (4)
    Let $G=A\times B$ for some groups $A$ and $B$. Suppose that $G$ acts coboundedly on a hyperbolic space $X$. Then the action of $A$ on $X$ is either elliptic or cobounded [1, Lemma 4.20].
  5. (5)
    Let $G$ be a group acting isometrically on a hyperbolic space $X$. Let $o\in X$ be a basepoint. Then an element $f\in G$ is a hyperbolic element if and only if the stable length $\tau(f):=\lim_{n\to\infty}\frac{d(o,f^{n}o)}{n}$ of $f$ is strictly bigger than 0.
  6. (6)
    Let $G$ be a group acting isometrically on a hyperbolic space $X$. Let $f\in G$ be a hyperbolic element. Then the subgroup $E(f):=\{g\in G:\exists\ n\geq 1,\text{ s.t. }gf^{n}g^{-1}=f^{n}\}$ admits a lineal action on $X$ with $\Lambda(E(f))=\Lambda\langle f\rangle$.
Hyperbolic structures
For a group $G$ with a generating set $S$, we denote by $\mathcal{G}(G,S)$ the Cayley graph of $G$ with respect to $S$.
Definition 5.2.
[1, Definition 1.1] Let $S$ and $T$ be two generating sets of a group $G$. We say that $S$ is dominated by $T$, written $S\preceq T$, if the identity map on $G$ induces a Lipschitz map between metric spaces $\mathcal{G}(G,T)\to\mathcal{G}(G,S)$. It is clear that $\preceq$ is a preorder on the set of generating sets of $G$ and therefore it induces an equivalence relation in the standard way:
S $S\text{ is equivalent to }T\Longleftrightarrow S\preceq T\text{ and }T\preceq$ S.
[…]
(4) $S]\preccurlyeq[T]\Longleftrightarrow S\preceq T.$
Definition 5.3.
[…]
Since hyperbolicity of a geodesic metric space is a quasi-isometry invariant, the definition above is independent of the choice of a particular representative in the equivalence class $[S]$.
Using the standard argument from the proof of the Svarc-Milnor lemma, it is easy to show that elements of $\mathcal{H}(G)$ are in one-to-one correspondence with equivalence classes of cobounded actions of $G$ on hyperbolic spaces considered up to a natural equivalence: two actions $G\curvearrowright X$ and $G\curvearrowright Y$ are equivalent if there is a coarsely $G$-equivariant quasi-isometry $X\to Y$. This gives another equivalent definition of property (PH’).
Definition 5.4 (Second definition of property (PH’)).
A (not necessarily finitely generated) group $G$ has property (PH’) if there exist finitely many hyperbolic structures $S_{1}],\cdots,[S_{n}]\in\mathcal{H}(G)$ such that the diagonal action of $G$ on $\prod_{i=1}^{n}\mathcal{G}(G,S_{i})$ equipped with $\ell^{1}$-metric is proper.
5.2. Inherited cobounded actions on hyperbolic spaces
In this subsection, we are going to show that cobounded group actions on hyperbolic spaces are inherited by normal quasi-retracts. At first, we need two simple lemmas.
Lemma 5.5.
Let $G$ be a group and $H$ be a normal subgroup of $G$. Then for any subset $T\subset G$ and $n\geq 1$, $\mathcal{C}(H,T^{n})\subset\mathcal{C}(H,T)^{n}$.
Proof.
For any $h\in H,g\in T^{n-1},t\in T$, observe that
$h,tg]=htgh^{-1}g^{-1}t^{-1}=h(gh^{-1}g^{-1}\cdot ghg^{-1})tgh^{-1}g^{-1}t^{-1}=[h,g][ghg^{-1},t].$
Then the conclusion follows from the above equality and mathematical inductions on $n$. ∎
Lemma 5.6.
Let $G$ be a group and $H$ be a normal subgroup of $G$. Let $T\subset G$ be a subset almost commuting with $H$. Then for any $h\in H,t\in T$ and $n\geq 1$, $(ht)^{n}\sim_{\mathcal{C}(H,T)^{n}}h^{n}t^{n}$.
Proof.
By Lemma 4.2, $\mathcal{C}(H,T)=\mathcal{C}(T^{-1},H)$. It suffices to show that $a_{n}=t^{-n}h^{-n}(ht)^{n}\in\mathcal{C}(T^{-1},H)^{n}$.
For each $1\leq k\leq n$, we denote $a_{n,k}:=t^{-(n-k)}h^{-(n-1)}(th)^{n-k}$. Since $H\lhd G$, it is not hard to see that $a_{n,k}\in H$ for each $1\leq k\leq n$. Observe that
$a_{n,k}=t^{-(n-k)}h^{-(n-1)}(th)^{n-k}=t^{-1}\cdot t^{-(n-k-1)}h^{-(n-1)}(th)^{n-k-1}\cdot t\cdot h=[t^{-1},a_{n,k+1}]a_{n,k+1}h$
for each $1\leq k\leq n-1$ and
$a_{n,n-1}h^{n-2}=t^{-1}h^{-(n-1)}th^{n-1}=[t^{-1},h^{-(n-1)}]=[t^{-1},a_{n,n}].$
Therefore,
$a_{n}$ $=t^{-n}h^{-n}(ht)^{n}=t^{-n}h^{-(n-1)}(th)^{n-1}t$
$=t^{-1}\cdot t^{-(n-1)}h^{-(n-1)}(th)^{n-1}\cdot t=[t^{-1},a_{n,1}]a_{n,1}$
$=[t^{-1},a_{n,1}][t^{-1},a_{n,2}]a_{n,2}h$
$=[t^{-1},a_{n,1}]\cdots[t^{-1},a_{n,n-1}]a_{n,n-1}h^{n-2}$
$=[t^{-1},a_{n,1}]\cdots[t^{-1},a_{n,n}]\in\mathcal{C}(T^{-1},H)^{n}$
and the conclusion follows. ∎
Let $G$ be a group and $H$ be a normal retract of $G$. Suppose $G$ admits a cobounded action on a hyperbolic space $X$. By Theorem 4.1, $G$ is isomorphic to $H\times G/H$. Then by Fact 1 (4), the action of $H$ on $X$ is either elliptic or cobounded. This implies that cobounded actions on hyperbolic spaces are inherited by normal retracts. Next, we are going to show the same conclusion holds for normal quasi-retracts.
Lemma 5.7.
Let $(G,H)$ be a group pair and $T$ be a right coset transversal of $H$ in $G$ which almost commutes with $H$. Suppose $G$ admits a cobounded action on a hyperbolic space $X$. If there exists a hyperbolic element in $H\cup T$, then the action $H\curvearrowright X$ is either elliptic or cobounded.
Proof.
For each hyperbolic element $f\in G$, we denote $E(f):=\{g\in G:\exists\ n\geq 1,\text{ s.t. }gf^{n}g^{-1}=f^{n}\}$. By assumption, we can pick a hyperbolic element $f\in H\cup T$.
Case I: $f\in T$.
Since $\mathcal{C}(H,T)$ is finite, one gets that $H\leq E(f)$. Since the action of $E(f)$ on $X$ is lineal, the action $H\curvearrowright X$ is either elliptic or lineal. If it is elliptic, then we are done. Now suppose the action $H\curvearrowright X$ is lineal. Let $h\in H$ be a hyperbolic element. Then $\Lambda(E(h))=\Lambda(H)=\Lambda(E(f))$. Similarly, since $\mathcal{C}(H,T)$ is finite, one gets that $T\subset E(h)$. Since $G=HT$, we conclude that $\Lambda(G)=\Lambda(H)$. This implies that the action of $G$ on $X$ is lineal. Hence, $X$ is a quasi-line and $H$ acts coboundedly on $X$.
Case II: $f\in H$.
Since $\mathcal{C}(H,T)$ is finite, one gets that $T\subset E(f)$. In particular, $\langle T\rangle\leq E(f)$. Since the action of $E(f)$ on $X$ is lineal, the action $\langle T\rangle\curvearrowright X$ is either elliptic or lineal. If it is elliptic, then the orbit of $G$ is within a finite Hausdorff distance to the orbit of $H$. Thus $H\curvearrowright X$ is cobounded since $G\curvearrowright X$ is cobounded. Now suppose the action $\langle T\rangle\curvearrowright X$ is lineal. Let $t\in\langle T\rangle$ be a hyperbolic element. Then $\Lambda(E(t))=\Lambda\langle T\rangle=\Lambda(E(f))$. Since $\mathcal{C}(H,T)$ is finite, it follows from Lemma 5.5 that $\mathcal{C}(H,t)$ is finite. Similarly, one gets that $H\leq E(t)$. Since $G=HT=H\langle T\rangle$, we conclude that $\Lambda(G)=\Lambda(E(t))$. This implies that the action of $G$ on $X$ is lineal. Hence, $X$ is a quasi-line and $H$ acts coboundedly on $X$.
In summary, we complete the proof. ∎
Lemma 5.8.
Let $G$ be a group and $H$ be a normal quasi-retract of $G$. Suppose $G$ admits a cobounded action on a hyperbolic space $X$. Then the action $H\curvearrowright X$ is either elliptic or cobounded.
Proof.
By Theorem 4.5, there exists a section $s:G/H\to G$ such that $s$ is a quasi-homomorphism and $\mathcal{C}(H,s(G/H))$ is finite. Denote $T=s(G/H)$ for simplicity. Note that any element $g\in G$ can be uniquely written as $g=g_{H}g_{T}$ where $g_{H}\in H$ and $g_{T}\in T$.
Since the action $G\curvearrowright X$ is cobounded, it belongs to one of the following types: elliptic, lineal, focal and of general type. If the action $G\curvearrowright X$ is elliptic, then $X$ is a bounded metric space and the conclusion holds obviously. If the action $G\curvearrowright X$ is lineal, then $X$ is a quasi-line and the conclusion follows from Fact 1 (2). Now suppose that the action $G\curvearrowright X$ is either focal or of general type.
Fix a basepoint $o\in X$. Denote $M:=\max_{g\in\mathcal{C}(H,T)}d(o,go)$. Pick a hyperbolic element $f\in G$ such that its stable length satisfies $\tau(f)>M$. By the above arguments, $f=f_{H}f_{T}$ where $f_{H}\in H$ and $f_{T}\in T$. By Lemma 5.6, for any $n\geq 1$, there exist $a_{1},\cdots,a_{n}\in\mathcal{C}(H,T)$ such that $f^{n}=(f_{H}f_{T})^{n}=f_{H}^{n}f_{T}^{n}a_{1}\cdots a_{n}$. By the triangle inequality,
$d(o,f^{n}o)=d(o,f_{H}^{n}f_{T}^{n}a_{1}\cdots a_{n}o)\leq d(o,f_{H}^{n}o)+d(o,f_{T}^{n}o)+\sum_{i=1}^{n}d(o,a_{i}o)\leq d(o,f_{H}^{n}o)+d(o,f_{T}^{n}o)+nM.$
By dividing by $n$ on both sides and letting $n\to\infty$, one gets that $0<\tau(f)-M\leq\tau(f_{H})+\tau(f_{T})$. This implies that at least one of $f_{H}$ and $f_{T}$ is hyperbolic. The remaining proof is completed by Lemma 5.7. ∎
Remark 5.9.
The conclusion of Lemma 5.8 does not hold for general quasi-retracts. For example, let $G=\mathbb{F}_{2}$ with a free basis $\{a,b\}$ and $H=\langle b\rangle$. Clearly $H$ is a retract of $G$. Define a homomorphism $\rho:G\to\mathrm{SL}_{2}(\mathbb{R})$ by mapping
$a=\left(\begin{array}[]{cc}\sqrt{2}&0\\ 0&1/\sqrt{2}\\ \end{array}\right)\text{ and }b=\left(\begin{array}[]{cc}1&1\\ 0&1\\ \end{array}\right).$
Then we obtain an action of $G$ on $\mathbb{H}^{2}$. This action is focal and cobounded (cf. [30, Proposition 4.5]). However, the orbit of $H$ in $\mathbb{H}^{2}$ is not even quasi-convex.
As a corollary, we have
Corollary 5.10.
Property (PH’) is inherited by normal quasi-retracts.
For an independent interest, we give a classification about the action of the defect subgroup with respect to a quasi-homomorphism on a hyperbolic space.
Lemma 5.11.
Let $\phi:G\to H$ be a surjective quasi-homomorphism with $D=D(\phi)$. Suppose $H$ admits a cobounded action on a hyperbolic space $X$. Then the action of the defect subgroup $\Delta_{\phi}$ on $X$ is either elliptic or lineal. Moreover, one of the following holds:
  1. (1)
    if $\Delta_{\phi}\curvearrowright X$ is lineal, then $\Delta_{\phi}\curvearrowright X$ is cobounded and $X$ is a quasi-line.
  2. (2)
    if $\Delta_{\phi}\curvearrowright X$ is elliptic, then the quotient space $X/\Delta_{\phi}$ endowed with the quotient metric is quasi-isometric to $X$ and thus is still a hyperbolic space.
Proof.
By Lemma 2.2, $\Delta_{\phi}=\langle D\rangle$ is a virtually abelian normal subgroup of $H$. Since any finitely generated group admitting a focal or general type action on a hyperbolic space has exponential growth, the action of $\Delta_{\phi}$ on $X$ is either elliptic, or horocyclic, or lineal.
Claim.
The action $\Delta_{\phi}\curvearrowright X$ is not horocyclic.
Proof of Claim.
Suppose to the contrary that the action $\Delta_{\phi}\curvearrowright X$ is horocyclic. Then every element of $\Delta_{\phi}$ is either elliptic or parabolic on $X$. A well-known fact is that if a finitely generated virtually abelian group acts isometrically on a hyperbolic space with each element being elliptic, then the action is also elliptic. Hence, there exists at least one parabolic element, say $f$, in $\Delta_{\phi}$ on $X$. Since $\Delta_{\phi}$ is finitely generated by $D$, there exists $N\geq 1$ such that $f\in D^{N}\cup D^{-N}$. By Lemma 2.1 (3), for any $h\in H$ and $n\geq 1$, $h^{n}fh^{-n}\in A:=(D^{2}D^{-1})^{N}\cup(D^{2}D^{-1})^{-N}$. Fix a basepoint $o\in X$ and denote $M:=\max_{g\in A}d(o,go)$.
Since $\Delta_{\phi}\lhd H$ and $H\curvearrowright X$ is cobounded, the action of $H$ on $X$ must be focal (cf. [1, Lemma 4.21]). Then there exists a hyperbolic element $h\in H$ such that $d(o,h^{n}fh^{-n}o)\to\infty$ as $n\to\infty$. This contradicts with $d(o,h^{n}fh^{-n}o)\leq M$ for any $n\geq 1$. Therefore, the action $\Delta_{\phi}\curvearrowright X$ is not horocyclic. ∎
By the above Claim, the action of $\Delta_{\phi}$ on $X$ is either elliptic or lineal.
If $\Delta_{\phi}\curvearrowright X$ is lineal, then it follows from $\Delta_{\phi}\lhd H$ that the action $H\curvearrowright X$ is also lineal. Since $H\curvearrowright X$ is also cobounded, one gets that $X$ is a quasi-line and thus $\Delta_{\phi}\curvearrowright X$ is also cobounded.
If $\Delta_{\phi}\curvearrowright X$ is elliptic, then the conclusion can be deduced from [3, Lemma 4.10]. ∎
5.3. Induced quasi-actions
Definition 5.12.
[21, Definition 2.3] A $(\lambda,\epsilon)$-quasi-action of a group $G$ on a metric space $X$ is a map $\rho:G\times X\to X$, denoted $\rho(g,x)\mapsto gx$, so that the following hold:
  • (i)
    for each $g$, $\rho(g,-):X\to X$ is a $(\lambda,\epsilon)$-quasi-isometry;
  • (ii)
    for each $x\in X$ and $g,h\in G$, we have
    $d(g(hx),(gh)x)=d(\rho(g,\rho(h,x)),\rho(gh,x))\leq\epsilon.$
(Note that $\lambda$ and $\epsilon$ must be independent of $g$ and $h$.) We call a quasi-action cobounded if, for every $x\in X$, the map $\rho(-,x):G\to X$ is $\epsilon^{\prime}$-coarsely onto for some $\epsilon^{\prime}\geq 0$.
Definition 5.13.
[21, Definition 2.4] Suppose that $\rho_{X}:G\times X\to X$ and $\rho_{Y}:G\times Y\to Y$ are quasi-actions. A map $\sigma:X\to Y$ is called coarsely equivariant if there is some $\epsilon\geq 0$ so that $d(\sigma\circ\rho_{X}(g,x),\rho_{Y}(g,\sigma(x)))\leq\epsilon$ for all $g\in G$ and $x\in X$. A coarsely equivariant quasi-isometry is called a quasi-conjugacy. If there exists a quasi-conjugacy between $X$ and $Y$, then the two quasi-actions are called quasi-conjugate.
By definition, two equivalent cobounded actions are quasi-conjugate.
Example 5.14.
Let $\phi:G\to\mathbb{R}$ be a quasimorphism with $\|D(\phi)\|_{\infty}=\epsilon$. A $(1,\epsilon)$-quasi-action of $G$ on $\mathbb{R}$ is given by $\rho(g,x)=x+\phi(g)$.
Let $\phi:G\to H$ be a quasi-homomorphism with $D=D(\phi)$. Suppose $H$ admits a $(\lambda,\epsilon)$-quasi-action on a metric space $X$ and there exists a constant $M>0$ such that $\sup_{x\in X,a\in D}d(x,ax)\leq M$. Define a map $\rho:G\times X\to X$ by $\rho(g,x):=\phi(g)x$. For any $g,h\in G$, there exists $a\in D$ such that $\phi(gh)=\phi(g)\phi(h)a$. Thus, for each $x\in X$ and $g,h\in G$, we have
$d(\rho(g,\rho(h,x)),\rho(gh,x))$ $=d(\phi(g)(\phi(h)x),\phi(gh)x)\leq d((\phi(g)\phi(h))x,(\phi(g)\phi(h)a)x)+\epsilon$
$\leq d((\phi(g)\phi(h))x,(\phi(g)\phi(h))(ax))+2\epsilon\leq\lambda d(x,ax)+3\epsilon\leq\lambda M+3\epsilon.$
The above inequality shows that $\rho$ is a $(\lambda,\lambda M+3\epsilon)$-quasi-action of $G$ on $X$. This quasi-action is referred to as an induced quasi-action by $\phi$. Furthermore, if $\phi$ is a quasi-retraction, then the induced quasi-action of $G$ on $X$ is an extension of the original quasi-action of $H$ on $X$.
Lemma 5.15.
Let $\phi:G\to H$ be a coarsely surjective quasi-homomorphism. Suppose $H$ admits a cobounded quasi-action on a metric space $X$. Then the map $\rho:G\times X\to X$ defined by $\rho(g,x):=\phi(g)x$ is an induced quasi-action of $G$ on $X$ by $\phi$.
Moreover, suppose $H$ admits another cobounded quasi-action on a metric space $Y$. Then the quasi-action of $H$ on $X$ is quasi-conjugate to the quasi-action of $H$ on $Y$ if and only if the induced quasi-action of $G$ on $X$ by $\phi$ is quasi-conjugate to the induced quasi-action of $G$ on $Y$ by $\phi$.
Proof.
Let $D=D(\phi)$ be the defect set of $\phi$. By the above arguments, it suffices to show that there exists a constant $M>0$ such that $\sup_{x\in X,a\in D}d(x,ax)\leq M$.
Since $\phi$ is coarsely surjective, there exists a constant $C>0$ such that $H=\mathcal{N}_{C}(\phi(G))$. Since $H$ is discrete, this is equivalent to saying that there exists a finite subset $A\subset H$ such that $H=\phi(G)\cdot A$. Fix a basepoint $o\in X$. Denote $M_{1}:=\max_{b\in A^{-1}D^{2}D^{-1}A}d(o,bo)$. Since the quasi-action of $H$ on $X$ is cobounded, there exists a constant $M_{2}>0$ such that for any $x\in X$, there exists $h_{x}\in H$ such that $d(x,h_{x}o)\leq M_{2}$. Since $H=\phi(G)A$, there exists $g_{x}\in G$ and $a_{x}\in A$ such that $h_{x}=\phi(g_{x})a_{x}$. Let $\lambda\geq 1,\epsilon\geq 0$ be two constants such that $H$ admits a $(\lambda,\epsilon)$-quasi-action on $X$. By the triangle inequality and Lemma 2.1 (3), we have that
$d(x,ax)$ $\leq d(x,h_{x}o)+d(h_{x}o,a(h_{x}o))+d(a(h_{x}o),ax)\leq d(h_{x}o,(ah_{x})o)+(\lambda+1)M_{2}+2\epsilon$
$\leq\lambda d(o,h_{x}^{-1}(ah_{x}o))+(\lambda+1)M_{2}+3\epsilon\leq\lambda d(o,(h_{x}^{-1}ah_{x})o)+(\lambda+1)M_{2}+4\epsilon$
$=\lambda d(o,(a_{x}^{-1}\phi(g_{x})^{-1}a\phi(g_{x})a_{x})o)+(\lambda+1)M_{2}+4\epsilon\leq\lambda M_{1}+(\lambda+1)M_{2}+4\epsilon.$
By setting $M=\lambda M_{1}+(\lambda+1)M_{2}+4\epsilon$, one gets that $\rho$ is a $(\lambda,\lambda M+3\epsilon)$-quasi-action of $G$ on $X$.
Moreover, suppose that there exists a coarsely $H$-equivariant quasi-isometry $\sigma:X\to Y$. By the above constructions, this map $\sigma$ gives also an induced coarsely $G$-equivariant quasi-isometry $\sigma:X\to Y$. And vice versa since $\phi$ is coarsely surjective. ∎
Definition 5.16.
[21, Definition 1.2] A group $G$ has property (QFA) if for every quasi-action of $G$ on any tree $X$, there is some $x\in X$ so that the orbit $Gx$ has finite diameter.
Lemma 5.17.
Let $G,H$ be two finitely generated groups and $\phi:G\to H$ be a coarsely surjective quasi-homomorphism. If $G$ has property (QFA), then so is $H$.
Proof.
Let $X$ be an arbitrary tree on which $H$ quasi-acts. By [21, Remark 3.2], the quasi-action of $H$ on $X$ can be assumed to be cobounded. As a result of Lemma 5.15, $\phi$ induces a cobounded quasi-action of $G$ on $X$. Since $G$ has property (QFA), there is some $x\in X$ so that the orbit $Gx:=\phi(G)x$ has finite diameter. Since $\phi$ is coarsely surjective, the orbit $Hx$ has also finite diameter. Therefore, $H$ has property (QFA) since $X$ is arbitrary. ∎
Corollary 5.18.
Property (QFA) is a quasi-isomorphism invariant for finitely generated groups.
Proposition 5.19.
Let $1\to H\to G\to Q\to 1$ be a short exact sequence of finitely generated groups which is left quasi-split. Then $G$ has property (QFA) if and only if both $H$ and $Q$ have property (QFA).
Proof.
[…]
To obtain the second item of Theorem 1.10, we need the following fact.
Proposition 5.20.
[20, Proposition 4.4] Any quasi-action on a geodesic metric space $X$ is quasi-conjugate to an action on some connected graph quasi-isometric to $X$. Conversely, any isometric action on a geodesic metric space quasi-isometric to $X$ is quasi-conjugate to some quasi-action on $X$.
Lemma 5.21.
Property (PH’) is a quasi-isomorphism invariant for discrete groups.
Proof.
Let $G,G^{\prime}$ be two discrete quasi-isomorphic groups. Suppose that $G^{\prime}$ has property (PH’) from actions $G^{\prime}\curvearrowright X_{1},\cdots,G^{\prime}\curvearrowright X_{n}$. Let $\phi:G\to G^{\prime}$ be a quasi-isomorphism. By Lemma 5.15, $\phi$ induces a cobounded quasi-action of $G$ on each $X_{i}$. As a result of Proposition 5.20, for each $1\leq i\leq n$, the induced quasi-action of $G$ on $X_{i}$ is quasi-conjugate to an action of $G$ on a hyperbolic space $X_{i}^{\prime}$. Via these quasi-conjugacy, we know that each action $G\curvearrowright X_{i}^{\prime}$ is also cobounded. Moreover, since $\phi$ is a quasi-isomorphism, the (PH’) property of $G^{\prime}$ implies that the diagonal action of $G$ on the $\ell^{1}$-product space $\prod_{i=1}^{n}X_{i}^{\prime}$ is proper. This gives the (PH’) property of $G$. ∎
Lemma 5.22.
[30, Lemma 7.1] Let $G$ be a group splitting as a direct product $G=H\times K$. Then
  1. (1)
    $G$ has property (PH’) if and only if both of $H,K$ have property (PH’).
  2. (2)
    $G$ has property (PH) if and only if both of $H,K$ have property (PH).
  3. (3)
    assume $G$ is finitely generated, $G$ has property (QT’) or (QT) if and only if both of $H,K$ have property (QT’) or (QT), respectively.
Proposition 5.23.
Let $1\to H\to G\to Q\to 1$ be a short exact sequence of discrete groups which is left quasi-split. Then $G$ has property (PH’) if and only if both $H$ and $Q$ have property (PH’).
Proof.
By Theorem 4.5, $G$ is strictly quasi-isomorphic to $H\times Q$. As a result of Lemma 5.21, $G$ has property (PH’) if and only if $H\times Q$ has property (PH’). Then the conclusion follows from Lemma 5.22 (1). ∎
As analogues of Lemma 5.21 and Proposition 5.23, we have
Lemma 5.24.
Property (QT’) is a quasi-isomorphism invariant for finitely generated groups.
Proof.
By [21, Remark 3.2], any isometric group action of a finitely generated group on a quasi-tree can be assumed to be cobounded. The remaining proof is similar to the proof of Lemma 5.21. ∎
In [7], Button showed that property (QT) is not a quasi-isometry invariant. Since quasi-isomorphisms are stronger equivalence relations than quasi-isometries for finitely generated groups, we can ask the following question.
Question 3.
Is property (QT) a quasi-isomorphism invariant for finitely generated groups?
By combining Lemmas 2.9, 5.21, 5.24 and Corollary 5.18, we obtain Proposition 1.9.
Proposition 5.25.
Let $1\to H\to G\to Q\to 1$ be a short exact sequence of finitely generated groups which is left quasi-split. Then $G$ has property (QT’) if and only if both $H$ and $Q$ have property (QT’).
Proof.
By Theorem 4.5, $G$ is strictly quasi-isomorphic to $H\times Q$. As a result of Lemma 5.24, $G$ has property (QT’) if and only if $H\times Q$ has property (QT’). Then the conclusion follows from Lemma 5.22 (3). ∎
To obtain the third item of Theorem 1.10, we need the following intermediate result.
In [26], Nguyen-Wang gave a geometric characterization for FZ groups via a notion of curvature for finitely generated groups. Here, we collect and give various equivalent algebraic characterizations for FZ groups, one of which is via quasi-retracts.
Proposition 5.26.
Let $G$ be a finitely generated group. The following are equivalent:
  1. (1)
    $G$ is FZ, i.e. $G:Z(G)]<\infty$.
  2. (2)
    $[G,G]$ is finite.
  3. (3)
    $\mathcal{C}(G,G)$ is finite.
  4. (4)
    $G$ is FC, i.e. each $g\in G$ has finitely many conjugates.
  5. (5)
    $G$ is finite-by-$\mathbb{Z}^{m}$.
  6. (6)
    $G$ is finite-by-abelian.
  7. (7)
    $G$ is virtually solvable and each of its subgroups is a quasi-retract.
  8. (8)
    $G$ has property (PH’) from orientable lineal structures, i.e. there exist finitely many quasi-lines $L_{1},\cdots,L_{n}$ on which $G$ acts orientable lineally such that the diagonal action of $G$ on $\prod_{i=1}^{n}L_{i}$ equipped with $\ell^{1}$-metric is proper.
Proof.
[…]
Denote $A_{0}=A$ and $A_{i}=[A_{i-1},A_{i-1}]$ for $i\geq 1$. Since $A$ is solvable, there exists $n\in\mathbb{N}$ such that $A_{n}=1$. Note that $A_{i}\lhd A_{i-1}\leq G$. By composing the embedding map $A_{i-1}\to G$ with the quasi-retraction $G\to A_{i}$, one gets that each $A_{i}$ is a quasi-retract of $A_{i-1}$ for $1\leq i\leq n$. As a result of Theorem 4.5, $A_{i-1}$ is strictly quasi-isomorphic to $A_{i}\times A_{i-1}/A_{i}$. Since strict quasi-isomorphism is an equivalent relation on discrete groups (cf. Lemma 2.8), one gets that $G$ is strictly quasi-isomorphic to $G/A\times\prod_{i=1}^{n}A_{i-1}/A_{i}$ which is almost commutative. By Lemma 2.9, $\mathcal{C}(G,G)$ is also finite. ∎
Remark 5.27.
Together with [30, Lemma 3.1], Proposition 5.26 (8) also gives the (QT’) property of finitely generated FZ groups.
As an immediate corollary of Lemma 2.9 and Proposition 5.26, we have
Corollary 5.28.
Being FZ is a quasi-isomorphism invariant for finitely generated groups.
Proposition 5.29.
Let $1\to H\to G\xrightarrow{\pi}Q\to 1$ be a short exact sequence of finitely generated groups which is left quasi-split. Suppose that $H$ is FZ. Then $G$ has property (PH) (resp. (QT)) if and only if $Q$ has property (PH) (resp. (QT)).
Proof.
At first, we show that “$G$ has property (PH) $\Leftrightarrow$ $Q$ has property (PH)”.
$\Rightarrow$”: Let $G^{\prime}$ be a finite-index subgroup of $G$ such that $G^{\prime}$ has property (PH’). Denote $H^{\prime}=H\cap G^{\prime}$ and $Q^{\prime}=\pi(G^{\prime})$. By the isomorphism theorem of groups, $G^{\prime}/(H\cap G^{\prime})\cong HG^{\prime}/H=\pi(G^{\prime})$. Thus, we obtain a new short exact sequence of groups as follows:
(5) $1\to H^{\prime}\to G^{\prime}\xrightarrow{\pi|_{G^{\prime}}}Q^{\prime}\to 1.$
Claim.
The short exact sequence (5) is left quasi-split.
Proof of Claim.
Since $H$ is FZ, Proposition 5.26 shows that $\mathcal{C}(H,H)$ is finite. Note that $H:H^{\prime}]\leq[G:G^{\prime}]<\infty$. As a result of Proposition 3.10, $H^{\prime}$ is a quasi-retract of $H$. Let $r_{1}:G\to H$ and $r_{2}:H\to H^{\prime}$ be two quasi-retractions. Then the composite map $r=r_{2}\circ r_{1}|_{G^{\prime}}:G^{\prime}\to H^{\prime}$ is a quasi-homomorphism and left-inverse to the inclusion map $H^{\prime}\to G^{\prime}$. This shows that $H^{\prime}$ is a quasi-retract of $G^{\prime}$ and the conclusion follows. ∎
By the above Claim, together with Proposition 5.23, one gets that $Q^{\prime}$ has property (PH’). Since $Q^{\prime}]\leq[G:G^{\prime}]<\infty, QQ$ has property (PH).
$\Leftarrow$”: Let $Q^{\prime}$ be a finite-index subgroup of $Q$ such that $Q^{\prime}$ has property (PH’). Denote $G^{\prime}=\pi^{-1}(Q^{\prime})$. Thus, we obtain a new short exact sequence of groups as follows:
$1\to H\to G^{\prime}\xrightarrow{\pi|_{G^{\prime}}}Q^{\prime}\to 1.$
By Proposition 5.26, $H$ has property (PH’). As a result of Proposition 5.23, $G^{\prime}$ has property (PH’). Since $G^{\prime}]=[Q:Q^{\prime}]<\infty, GG$ has property (PH).
By substituting Proposition 5.25 for Proposition 5.23, the same arguments show that $G$ has property (QT) if and only if $Q$ has property (QT). ∎
By combining Propositions 5.19, 5.23, 5.25 and 5.29, we obtain Theorem 1.10. Note that Proposition 5.29 requires an additional assumption on $H$. If the answer to the following question is yes, then we can remove this assumption.
Question 4.
Let $G$ be a group and $G^{\prime}\leq G$ be a finite-index subgroup. Suppose that $H$ is a normal quasi-retract of $G$. Is $H\cap G^{\prime}$ necessarily a normal quasi-retract of $G^{\prime}$?
Remark 5.30.
If one drops the word “normal” in the above question, then the answer is negative. In [28, Theorem A], Snopce-Tanushevski-Zalesskii proved that for a free group $\mathbb{F}_{n}$, there exists a finite-index subgroup $H$ and a retract $R$ of $\mathbb{F}_{n}$ such that $H\cap R$ is not a retract of $H$.
In the end of this subsection, we prove Proposition 1.11. The proof will be divided into the next two propositions.
Proposition 5.31.
Let $\phi:G\to H$ be a coarsely surjective quasi-homomorphism. Then $\mathcal{H}(H)$ embeds in $\mathcal{H}(G)$ as a sub-poset.
Proof.
For each $T]\in\mathcal{H}(H)$, $\mathcal{G}(H,T)$ is a hyperbolic graph on which $H$ acts coboundedly. By Lemma 5.15, $\phi$ induces a cobounded quasi-action of $G$ on $\mathcal{G}(H,T)$. As a result of Proposition 5.20, the induced cobounded quasi-action of $G$ on $\mathcal{G}(H,T)$ is quasi-conjugate to a cobounded action of $G$ on a hyperbolic graph $X_{T}$. By the classical Milnor-Svarc Lemma, there exists a generating set $S$ of $G$ such that $G\curvearrowright\mathcal{G}(G,S)$ is quasi-conjugate to $G\curvearrowright X_{T}$. Then we can define a map $\chi:\mathcal{H}(H)\to\mathcal{H}(G)$ by $\chi([T])=[S]$. It follows from the “Moreover” part of Lemma 5.15 that $\chi$ is well-defined and injective.
It remains to show that $\chi$ is order-preserving. Let $T]\preccurlyeq[T^{\prime}]$ be two hyperbolic structures in $\mathcal{H}(H)$. By definition, the identity map induces a Lipschitz map $\sigma:\mathcal{G}(H,T^{\prime})\to\mathcal{G}(H,T)$. According to Lemma 5.15, $\sigma$ is also an induced coarsely $G$-equivariant Lipschitz map from $\mathcal{G}(H,T^{\prime})$ to $\mathcal{G}(H,T)$. After applying Proposition 5.20 and the classical Milnor-Svarc Lemma, we get that there exists a coarsely $G$-equivariant coarsely Lipschitz map from $\mathcal{G}(G,S^{\prime})$ to $\mathcal{G}(G,S)$ where $S\in\chi([T])$ and $S^{\prime}\in\chi([T^{\prime}])$. This shows that $\chi([T])\preccurlyeq\chi([T^{\prime}])$.
In summary, $\chi$ is an order-preserving injection from $\mathcal{H}(H)$ to $\mathcal{H}(G)$. ∎
Proposition 5.32.
Let $G,G^{\prime}$ be two discrete quasi-isomorphic groups. Then $\mathcal{H}(G)\cong\mathcal{H}(G^{\prime})$.
Proof.
For each $S^{\prime}]\in\mathcal{H}(G^{\prime})$, $\mathcal{G}(G^{\prime},S^{\prime})$ is a hyperbolic graph on which $G^{\prime}$ acts coboundedly. Let $\phi:G\to G^{\prime}$ be a quasi-isomorphism. By Lemma 5.15, $\phi$ induces a cobounded quasi-action of $G$ on $\mathcal{G}(G^{\prime},S^{\prime})$. As a result of Proposition 5.20, the induced cobounded quasi-action of $G$ on $\mathcal{G}(G^{\prime},S^{\prime})$ is quasi-conjugate to a cobounded action of $G$ on a hyperbolic graph $X_{S^{\prime}}$. By the classical Milnor-Svarc Lemma, there exists a generating set $S$ of $G$ such that $G\curvearrowright\mathcal{G}(G,S)$ is quasi-conjugate to $G\curvearrowright X_{S^{\prime}}$. Then we can define a map $\chi:\mathcal{H}(G^{\prime})\to\mathcal{H}(G)$ by $\chi([S^{\prime}])=[S]$. It follows from the “Moreover” part of Lemma 5.15 that $\chi$ is well-defined and injective.
Since $\phi:G\to G^{\prime}$ is a quasi-isomorphism, there exists a quasi-inverse $\phi^{\prime}:G^{\prime}\to G$ such that $\mathrm{dist}(\phi^{\prime}\circ\phi,Id_{G})<\infty$ and $\mathrm{dist}(\phi\circ\phi^{\prime},Id_{G^{\prime}})<\infty$. Similarly, we can define a map $\chi^{\prime}:\mathcal{H}(G)\to\mathcal{H}(G^{\prime})$. It follows from the above definitions that $\chi\circ\chi^{\prime}([S])=[S]$ for any $S]\in\mathcal{H}(G)$. This shows that $\chi$ is surjective.
The property that $\chi$ is order-preserving follows from the same arguments in the proof of Proposition 5.31. In summary, $\chi$ is an order-preserving bijection from $\mathcal{H}(G^{\prime})$ to $\mathcal{H}(G)$. ∎
By combining Proposition 5.31 and Proposition 5.32, we complete the proof of Proposition 1.11.